### v1

# Two statistical problems for multivariate mixture distributions

Abstract.
In an earlier work [12], it was shown that mixtures of multivariate Gaussian or $t$-distributions can be distinguished by projecting them onto a certain predetermined finite set of lines, the number of lines depending only on the total number of distributions involved and on the ambient dimension. Using this work, we address the following two important statistical problems: that of testing and measuring the agreement between two different random partitions, and that of estimating for mixtures of multivariate normal distributions and mixtures of $t$-distributions based of univariate projections. We also compare our proposal with robust versions of the expectation-maximization method EM. In each case, we present algorithms for effecting the task, and compare them with existing methods by carrying out some simulations.
Key words and phrases:
clustering, Gaussian distribution, identifiability, mixture, projection, random partition, $t$-distribution
1. Introduction
Some recent results refine the classical Cramér–Wold theorem, and show that in some cases just finitely many projections suffice to determine the measure. See for instance [11] and the references therein, which include the case of Gaussian measures where just $(d^{2}+d)/2$ projections suffice.
Recently, in [12] these results have been extended to the case of mixtures, namely convex combinations of probability measures. In particular, the important cases of multivariate Gaussian and $t$-distributions mixtures are addressed in [12].
With these results in hand we then consider a new approach to two interesting statistical problems. The first one is a way to measure the agreement between two partitions of a data set, in particular, two outputs of different clustering procedures. The second one is to provide a new procedure for estimating a mixture of multivariate normal and $t$-distributions based on univariate projections.
Section 2 begins with a short introduction to the notion of the Cramér–Wold device. In Section 3 we discuss the problem of comparing random partitions by model-based methods. Comparing partitions is an important problem in particular in cluster analysis, and a large amount of work on the subject introduced several different proposals. The most popular measures are given by the Rand index [26], the adjusted Rand index (Hubert and Arabie [18]), the Jaccard index, the Dunn index, the silhouette index, and the Xie–Beni index among others. See for instance the reviews by Arabie and Boorman [1], Fowlkes and Mallows [10] among others. However, most of these existing indices lack real mathematical analysis, and almost no information exists about their distribution, see [20].
The way that we approach this problem is to measure the agreement between the probability measures that have produced those partitions (the output of two different model-based clustering procedures). Model-based approach provides a good interpretation of this measure in terms of the underlying unknown distributions. Indeed, if the mixtures distributions are close, then we can conclude that the partitions are close, even when they have been produced by different data bases, providing a more general setup to the problem.
Two algorithms implementing these ideas are described, and then illustrated with some simulations. The simulations indicate a high degree of agreement between several well-known distance-measures and the one that we propose.
Finally, in Section 4, we present a new procedure for estimating mixtures of multivariate normal distributions, and mixtures of $t$-distributions based on one-dimensional projections of the mixture based on characteristic functions (ECF (see [33] and [37]), and using those univariate mixtures to provide consistent estimators of the parameters of the multivariate mixture.
The proposed random-projection method (RP) is compared with the well-known expectation-maximization method (EM) and robust versions of EM in some simulations. We use some results in [33] (see also [39] and [37]) for some asymptotic properties. An important fact, mentioned in [33], is that “It is shown that, using Monte Carlo simulation, the finite sample properties of the ECF estimator are very good, even in the case where the popular maximum likelihood estimator fails to exist…”. We conclude with a further example using real data from the National Institute for Education Evaluation (INEEd) in Uruguay. The dataset is open access, and the codes, written in R, are available on GitHub [^1] https://github.com/mrleomr/MultivariateMixture or can be requested directly from the authors.
2. Some preliminaries and recent results to be used
In this section, we introduce the fundamental concepts necessary for understanding our work. In particular, we clarify what we mean by projecting a measure onto a subspace and how distributions can be constructed through mixtures of Gaussian or Student distributions.
2.1. Projections
Given a Borel probability measure $P$ on $\mathbb{R}^{d}$ and a vector subspace $H$ of $\mathbb{R}^{d}$, we denote the projection of $P$ onto $H$ as $P_{H}$, specifically defined as the Borel probability measure on $H$ given by
$P_{H}(B):=P(\pi_{H}^{-1}(B)),$
where $\pi_{H}:\mathbb{R}^{d}\to H$ is the orthogonal projection of $\mathbb{R}^{d}$ onto $H$.
Let $\mathcal{P}$ denote a family of Borel probability measures on $\mathbb{R}^{d}$, and let $\mathcal{H}$ represent a family of vector subspaces of $\mathbb{R}^{d}$ (not necessarily of the same dimension). We say that $\mathcal{H}$ constitutes a Cramér–Wold system for $\mathcal{P}$ if, for every pair $P,Q\in\mathcal{P}$,
$P_{H}=Q_{H}\leavevmode\nobreak\ (\forall H\in\mathcal{H})\quad\Rightarrow\quad P=Q.$
In this context, the original Cramér–Wold theorem states that the family of all lines in $\mathbb{R}^{d}$ constitutes a Cramér–Wold system for the family of Borel probability measures on $\mathbb{R}^{d}$.
2.2. Mixtures
Let $d\geq 1$ and let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$. A $\mathcal{P}$-mixture is defined as a convex combination of measures from $\mathcal{P}$. In other words, it takes the form
$\sum_{j=1}^{n}\lambda_{j}P_{j}$
where $P_{1},\dots,P_{n}\in\mathcal{P}$ and $\lambda_{1},\dots,\lambda_{n}\geq 0$ with $\sum_{j=1}^{n}\lambda_{j}=1$.
Once the data has been projected, we will focus on univariate distributions. Specifically, if $d=1$, we will consider the two families of univariate distributions: Gaussian distributions and $t$-distributions. Gaussian distributions are characterized by their densities
(1) $f_{\mu,\sigma}(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\Bigl{(}-\frac{(x-\mu)^{2}}{2\sigma^{2}}\Bigr{)}\quad(x\in\mathbb{R}).$
A $t$-distribution on $\mathbb{R}$ with $\nu$ degrees of freedom is a Borel measure characterized by the density function
$f_{\nu,\mu,\sigma}(x)=c_{\nu,\mu,\sigma}\left(1+\frac{(x-\mu)^{2}}{\nu\sigma^{2}}\right)^{-(\nu+1)/2},$
where $\nu$ is a positive integer, $\mu\in\mathbb{R}$ and $\sigma>0$. The constant $c_{\nu,\mu,\sigma}$ is chosen to ensure that
[…]
This distribution has a mean of $\mu$ (provided that $\nu>1$) and a variance of $\sigma^{2}\nu/(\nu-2)$ (if $\nu>2$).
In the multivariate case, where $d>1$, a Gaussian measure $P$ on $\mathbb{R}^{d}$ is defined by a density of the form
(2) $\frac{1}{(2\pi\det(\Sigma))^{d/2}}\exp\Bigl{(}-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\Bigr{)}\quad(x\in\mathbb{R}^{d}),$
where $\mu\in\mathbb{R}^{d}$, and where $\Sigma$ is a real $d\times d$ positive-definite matrix.
A Gaussian mixture is a measure on $\mathbb{R}^{d}$ that represents a finite convex combination of Gaussian measures. Mixtures of multivariate Gaussian distributions exhibit several advantageous properties. In particular, Titterington et al. [32] demonstrate that Gaussian kernel density estimators can approximate any continuous density given a sufficient number of kernels, establishing their universal consistency (see also Scott [30]). Additionally, it is well known that Gaussian mixtures are weak*-dense in the space of all Borel probability measures on $\mathbb{R}^{d}$.
Gaussian-mixture models have proven to be highly effective in modeling various real-world data. For a comprehensive examination of their properties, we refer to Titterington et al. [32]. These models are flexible and general, with relevant applications documented in numerous fields, including density estimation, machine learning, and clustering. As noted in [3], significant applications include Pearson’s work on modeling crab data [23], Roeder’s studies [28], and the research by Schork, Allison and Thiel in genetics [29], as well as Everitt’s investigations involving schizophrenia patients [9], among many others.
Estimating these models can be quite complex, particularly for high-dimensional data, and typically involves Markov-chain Monte-Carlo methods within a Bayesian framework. Important applications for cluster analysis are discussed in Tadesse, Sha, and Vannucci [31], as well as in Raftery and Dean [25].
Similarly, when the data exhibit heavy tails, a similar development is plausible using mixtures of Student distributions. A Student distribution ($t$-distribution) on $\mathbb{R}^{d}$ is a measure with density of the form
$f_{\nu,\mu,\Sigma}(x)=c_{\nu,\mu,\Sigma}\Bigl{(}1+\frac{(x-\mu)^{T}\Sigma^{-1}(x-\mu)}{\nu}\Bigr{)}^{-(\nu+d)/2}\quad(x\in\mathbb{R}^{d}),$
[…]
2.3. A Cramér–Wold theorem for Gaussian mixtures and $t$-mixtures
In this section, we address the problem of testing for equality between two Gaussian mixtures by analyzing a finite number of projections. The basic theorem supporting this approach consists of two key components. The first is the abstract result presented in Theorem 4.1 in [12]. The second is a characterization of Cramér–Wold systems for Gaussian measures in $\mathbb{R}^{d}$ (and, more generally, for elliptical distributions), established in [11, Theorems 1 and 2], which we will now summarize.
Let $S$ be a set of vectors in $\mathbb{R}^{d}$. The corresponding set of lines $\{\langle x\rangle:x\in S\}$ forms a Cramér–Wold system for the Gaussian measures in $\mathbb{R}^{d}$ if and only if $S$ satisfies the property that the only real symmetric $d\times d$ matrix $A$ for which $x^{T}Ax=0$ for all $x\in S$ is the zero matrix. A set $S$ with this property is referred to as a symmetric-matrix uniqueness set (or sm-uniqueness set for short).
In [11], it was demonstrated that an sm-uniqueness set for $\mathbb{R}^{d}$ must span $\mathbb{R}^{d}$and contain at least $(d^{2}+d)/2$ vectors. We will denote $S$ as a strong sm-uniqueness set if every subset of $S$ containing $(d^{2}+d)/2$ vectors is also an sm-uniqueness set.
We now recall the Cramér–Wold theorem for Gaussian mixtures [12, Theorem 4.1].
Theorem 2.1.
Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ Gaussian measures on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(\ell+m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
We also recall an analogue of Theorem 2.1 for mixtures of multivariate $t$-distributions, given in [12, Theorem 4.2], thereby allowing heavy-tailed distributions.
Theorem 2.2.
Let $P$ and $Q$ be convex combinations of $\ell$ and $m$ multivariate $t$-distributions on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(\ell+m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
2.4. Strong sm-uniqueness sets
Theorems 2.1 and 2.2 beg the question as to whether there exist strong sm-uniqueness sets of arbitrarily large cardinality. An affirmative answer is given in the following result based on [12, Theorem 4.3], which also provides a realistic method for generating them.
Theorem 2.3.
Let $d\geq 2$, let $k\geq(d^{2}+d)/2$, and let $v_{1},\dots,v_{k}$ be independent random vectors in $\mathbb{R}^{d}$ whose distributions are given by densities on $\mathbb{R}^{d}$. Then, with probability one, the set $\{v_{1},\dots,v_{k}\}$ is a strong sm-uniqueness set for $\mathbb{R}^{d}$.
This theorem allows us to easily generate strong sm-uniqueness sets, by uniformly sampling $k$ random directions on the unit sphere. With these results in hand we are ready to study the two statistical problems mentioned above.
3. Comparing random partitions by model-based methods
3.1. Introduction
The problem of comparing two different partitions of a finite set of observations is an important classical problem in the clustering literature. There is a large amount of work on the subject with several different proposals.
The most popular measures are given by the Rand index [26], the adjusted Rand index (Hubert and Arabie [18]), the Jaccard index, the Dunn index, the silhouette index, and the Xie–Beni index among others. See for instance the reviews by Arabie and Boorman [1], Fowlkes and Mallows [10] among others. More recently Youness and Saporta [38] proposed a different approach to the problem. Most of these existing indices lack real mathematical analysis, and almost no information exists about their distribution, see [20]. Also, the problem becomes harder for high-dimensional data. Our model-based approach provides a good interpretation of this measure in terms of the underlying unknown distributions.
On the other hand, a probabilistic model used quite often in Bayesian statistics as well in cluster analysis is to consider that the data are generated by a mixture of multivariate normal distributions. Indeed, a popular and well-studied clustering method is to use Gaussian mixture models (GMM) for data clustering (model-based clustering). One can perform hard clustering or soft clustering. For hard clustering the GMM model assigns each datum to the component that maximizes the component posterior probability, given the datum, and each datum is assigned to only one cluster. Soft clustering assigns each instance a probability of belonging to a cluster. See for instance Gormley et al. [14] for details.
From the recent article [14] we quote: “Through its basis in a statistical modeling framework, model-based clustering provides a principled and reproducible approach to clustering. In contrast to heuristic approaches, model-based clustering allows for robust approaches to parameter estimation and objective inference on the number of clusters, while providing a clustering solution that accounts for uncertainty in cluster membership …”
Instead of measuring agreement between two different partitions, we consider the problem of measuring the agreement between the probability measures that have produced those partitions (the output of two different model-based clustering procedures), when both distributions correspond to finite mixtures of multivariate normal or $t$-distributions. In particular, we are able to treat high-dimensional data, and to perform a bootstrap test for $H0$ being that both partitions have been generated by the same mixture distribution against $HA$ that the partitions have been generated by two different mixture distributions.
As pointed in Baudry et al. [2], several problems appear when performing Gaussian-model-based clustering, affecting the number of clusters obtained by BIC. Some nice solutions are proposed. However these problems do not affect us. An important aspect of our approach is that we do not need to perform the clustering method to calculate the measuring agreement, and that also we can easily deal with high-dimensional data. In other words, the method does not require a prior criterion for selecting the number of clusters, and therefore avoids the issues associated with poor estimation of the number of clusters.
3.2. Description of the algorithms
3.2.1. Two independent samples
Suppose that we want to compare the results obtained from two clustering procedures but from the same variables. We have two independent samples $\aleph_{1}=\{X_{1},\dots X_{\ell}\}\subset\mathbb{R}^{d}$ and $\aleph_{2}=\{Y_{1},\dots,Y_{m}\}\subset\mathbb{R}^{d}$, where $\aleph_{1}$ and $\aleph_{2}$ are iid samples with distributions $P,Q$ respectively, both being mixtures of multivariate normal or $t$-distributions.
We proceed as follows. Let $P_{\ell}$ and $Q_{m}$ be the empirical distributions of $\aleph_{1}$ and $\aleph_{2}$. Then we apply Theorem 2.1 (or Theorem 2.2 respectively), and report as a measure of discrepancy between the clusters the average of the projected Kolmogorov–Smirnov distance between the empirical distributions $P_{\ell}$ and $Q_{m}$, along the set of random directions given in Theorem 2.3.
More precisely, we perform the following algorithm.
Algorithm 3.1.
Fix an integer $k\geq\frac{1}{2}(l+m-1)(d^{2}+d-2)-1$.
(i) Take $k$ independent random directions $u_{1},\dots,u_{k}$, uniformly distributed in the unit sphere of $\mathbb{R}^{d}$. According to Theorem 2.3, the set $\{u_{1},\ldots,u_{k}\}$ is a strong sm-uniqueness set for $\mathbb{R}^{d}$ with probability one.
(ii) For each direction $u_{i}$, project the data $\aleph_{1}$ and $\aleph_{2}$ onto the direction $u_{i}$, and thereby obtain two univariate samples
$\mathcal{M}_{1i}$ =$=\{Z_{1i}=\langle u_{i},X_{1}\rangle,\dots,Z_{\ell i}=\langle u_{i},X_{\ell}\rangle\},$
$\mathcal{M}_{2i}$ =$=\{W_{1i}=\langle u_{i},Y_{1}\rangle,\ldots,W_{mi}=\langle u_{i},Y_{m}\rangle\}.$
(iv) For each $i\in\{1,\dots k\}$, calculate the Kolmogorov–Smirnov distance between the empirical distributions associated with the samples $\mathcal{M}_{1i}$ and $\mathcal{M}_{2i}$. Denote it by $\textit{KS}(i)$.
(v) Output: report as a measure of agreement $\textit{MA}_{k}:=\frac{1}{k}\sum_{i=1}^{k}\textit{KS}(i)$.
Remark.
Algorithm 3.1 can also be used for the two-sample testing problem for mixtures of Gaussian or $t$-distributions. Indeed, given the two samples $\aleph_{1}$ and $\aleph_{2}$ we consider the statistical test problem: $H0:P=Q$ vs $HA:P\neq Q$. Next we perform the algorithm and calculated $\textit{MA}_{k}$. Once $\textit{MA}_{k}$ has been calculated, the critical value can be found using bootstrap as in [11].
3.2.2. One-sample and different clustering procedures
Suppose now that we are given one sample =$\aleph=\{X_{1},\dots,X_{N}\}\subset\mathbb{R}^{d}$. We split the sample into two disjoint subsets $\aleph_{1},\aleph_{2}$. We apply to each subsample a different clustering procedure. For instance, we apply GMM with different prior distributions.
A first naive approach to the problem could be as follows. For each of the the clusters outputs, we want to define a distribution that corresponds to a Gaussian-mixture model. We build it up by considering the mean and the covariance matrix of each cluster, and the weights given by the proportion of points in each cluster. Let $\widehat{P}_{1},\widehat{P}_{2}$ denote those distributions. Robust estimators of the means and covariance matrices can be considered also. Next we generate two samples of the same size, one with distribution $\widehat{P}_{1}$ and the other one with distribution $\widehat{P}_{2}$. Finally we apply Algorithm 3.1 to these samples.
However this approach is too complicated, in particular for high-dimensional data. Since we will end up working with one-dimensional projections, we propose instead the following simpler algorithm. First we observe that, if the data is projected using a linear transformation and a GMM is fitted on the original data, then the density on the projection subspace is again a mixture of Gaussian variables with covariance matrices and mean vectors obtained by a linear transformation of the parameters estimated in the original scale of the data. Thus, for each random direction, we can fit a univariate Gaussian mixture of the projected data. We estimate for each direction the mean and the variance of the projected data, and we use the weights given by the proportion of points at each cluster.
More precisely, we perform the following algorithm:
Algorithm 3.2.
We start with two partitions outputs of the clustering procedures: $\{C_{1},\dots,C_{\ell}\}$ and $\{E_{1},\ldots,E_{m}\}$.
(i) Let $\alpha_{1},\ldots,\alpha_{\ell}$ the proportion of points in the sample at each cluster with procedure $1$, and $\beta_{1},\ldots,\beta_{m}$ the same for procedure $2$.
(ii) For each random direction $u_{i}\leavevmode\nobreak\ (i=1\dots k)$, we consider the projected samples of $C_{j}\leavevmode\nobreak\ (j=1,\dots\ell)$ on the direction $u_{i}$, and obtain estimators of their mean and variance, $\widehat{\mu}_{1ij}$ and $\widehat{\sigma^{2}}_{1ij}$.
(iii) For each random direction $u_{i}\leavevmode\nobreak\ (i=1,\ldots k)$, we consider the projected samples of $E_{j}\leavevmode\nobreak\ (j=1,\dots m)$ on the direction $u_{i}$, and obtain estimators of their mean and variance $\widehat{\mu}_{2ij}$ and $\widehat{\sigma^{2}}_{2ij}$.
(iv) Now consider, for each direction $u_{i}$, the one-dimensional Gaussian mixtures given by
$\widehat{F}_{1i}:=\sum_{j=1}^{\ell}\alpha_{j}N(\widehat{\mu}_{1ij},\widehat{\sigma}^{2}_{1ij})\quad\text{and}\quad\widehat{F}_{2i}:=\sum_{j=1}^{m}\beta_{j}N(\widehat{\mu}_{2ij},\widehat{\sigma}^{2}_{2ij}).$
Using Monte-Carlo, calculate $\textit{KS}(i):=\textit{KS}(\widehat{F}_{1i},\widehat{F}_{2i})$.
(v) Output: report as a measure of agreement $\textit{MA}_{k}:=\frac{1}{k}\sum_{i=1}^{k}\textit{KS}(i)$.
The measure $\textit{MA}_{k}$ takes values between 0 and 1. If the distributions
${F}_{1i}:=\sum_{j=1}^{\ell}\alpha_{j}N(\mu_{1ij},\sigma^{2}_{1ij})\quad\text{and}\quad F_{2i}:=\sum_{j=1}^{m}\beta_{j}N(\mu_{2ij},\sigma^{2}_{2ij}),$
are equal, then $\textit{MA}_{k}=0$, and varies according to the difference between the distribution of the mixtures fitted by both partitions, increasing as long as they are further away.
3.3. Simulations
3.3.1. A simple example of distances between mixtures
Consider two mixtures of two bivariate Gaussians, with parameters
$F_{1}$ $:=\lambda_{1}N\left(\mu_{1,1},\Sigma_{1,1}\right)+(1-\lambda_{1})N\left(\mu_{2,1},\Sigma_{2,1}\right),$
$F_{2}$ $:=\lambda_{2}N\left(\mu_{1,2},\Sigma_{1,2}\right)+(1-\lambda_{2})N\left(\mu_{2,2},\Sigma_{2,2}\right),$
where
$\mu_{1,1}$ $=\mu_{1,2}:=(1,-1),\mu_{2,1}:=(-2,2),\lambda_{1}=0.5,$
$\Sigma_{1,1}$ $=\Sigma_{1,2}:=\begin{pmatrix}1&0\\ 0&2\end{pmatrix}\quad\text{and}\quad\Sigma_{2,1}:=\begin{pmatrix}3&1\\ 1&4\end{pmatrix},$
and where, given $\eta_{1},\eta_{2},\eta_{3}$, we set $\mu_{2,2}:=\mu_{2,1}+(\eta_{1},\eta_{1})$, $\Sigma_{2,2}:=(1+\eta_{2})\Sigma_{2,1}$, and $\lambda_{2}:=0.5+\eta_{3}$.
The distances between the two mixtures $F_{1}$ and $F_{2}$ are calculated for three scenarios:
Scenario 1:
$\eta_{1}$ on a grid of values in $[0,1]$ and $\eta_{2}=\eta_{3}=0$,
Scenario 2:
$\eta_{2}$ on a grid of values in $[0,2]$ and $\eta_{1}=\eta_{3}=0$,
Scenario 3:
$\eta_{3}$ on a grid of values in $[0,0.5]$ and $\eta_{1}=\eta_{2}=0$.
Figure 1. Simulation per Scenario is provided considering $\eta_{1}=1$, $\eta_{2}=2$, and $\eta_{3}=1/2$ in Scenarios 1, 2, and 3, respectively
(a) Scenario 1
(b) Scenario 2
(c) Scenario 3
Figure 2. Boxplot of $500$ replicates of $\textit{MA}_{k}$ calculated between mixtures $F_{1}$ and $F_{2}$ by varying $\eta_{1},\eta_{2}$ and $\eta_{3}$ in the scenarios.
In Figure  1, for better visualization, a simulation per scenario is provided considering $\eta_{1}=1$, $\eta_{2}=2$, and $\eta_{3}=1/2$ in Scenarios 1, 2, and 3, respectively. Figure 2 shows how, when the two mixtures $F_{1}$ and $F_{2}$ diverge in location averages, dispersion or weights of the mixture components respectively, the $\textit{MA}_{k}$ average proposed in Algorithm 3.2 captures this separation.
For this purpose, in each scenario, for the selected values of $\eta_{1}$,$\eta_{2}$ and $\eta_{3}$, $500$ values of each mixture are simulated and $\textit{MA}_{k}$ is calculated. This procedure is replicated $500$ times, the Figure 2 shows the boxplots of these $500$ replicates for each case.
When all three parameters, $\eta_{1}$,$\eta_{2}$ and $\eta_{3}$, are zero, both mixtures coincide and the distance between the data simulated by both should be close. As the mixtures are separated, the $\textit{MA}_{k}$ values show an increasing behaviour.
3.3.2. An example of distances between mixtures for quality clustering
An interesting application of Algorithm  3.2 is to measure the quality of a clustering procedure, as discussed previously. In this example we illustrate how the classification error is captured by the distance between mixtures.
Assume that the cluster number $\ell$ is known a priori. We define $\psi_{i}=j$ if the true label of observation $X_{i}$ is $C_{j}$ with $i=1,2,\ldots,n$ and $j=1,2,\ldots,\ell$. Let $\widehat{\psi}_{i}$ be the labels assigned by the clustering procedure, and denote by $\Pi_{\ell}$ the family of permutations of $\{1,2,\ldots,\ell\}$.
The prediction error is given by the minimum between all permutations of the assigned labels and the true labels, that is,
$\min_{\pi\in\Pi_{\ell}}\frac{1}{k}\sum_{i=1}^{k}\mathbbm{1}_{\left\{\psi_{i}\neq\pi\left(\widehat{\psi}_{i}\right)\right\}}.$
On the other hand, $\textit{MA}_{k}$ is calculated between the true mixture and the one generated by the clustering procedure, where $\widehat{\lambda}_{j}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}_{\left\{\widehat{\psi}_{i}=j\right\}}$.
We consider $\ell=2$ and generate a sample of size $n=500$ from the mixture of two bivariate Gaussians distributions, with parameters
$\lambda$ $:=(0.3,0.7),\ \mu_{1}:=(1,-1)\ \quad\text{and}\quad\mu_{2}=(-2,2),$
$\Sigma_{1}$ $:=\begin{pmatrix}1&0.5\\ 0.5&2\end{pmatrix}\quad\text{and}\quad\Sigma_{2}:=\begin{pmatrix}3&1\\ 1&4\end{pmatrix}.$
The clustering methods implemented are $k$-means, see [19], $k$-medoids (PAM), [27], Spectral clustering, see [34], Hierarchical Ward, see [35] and Mean Shift, [4]. These algorithms were implemented in the R language with the packages stats, cluster, hclust, kernlab and LPCM.
[…]
Figure 3. Estimation of the deviations of the distances obtained for the two mixtures in Scenario 1 when $\eta=1$, as the number of projections $k$ increases.
[…]
4. A new estimator for Gaussian mixtures
4.1. Introduction
In this section, we shall develop a new estimator of a multivariate mixture of Gaussian distributions based on two ideas: random projections to obtain one-dimensional mixtures of normal distributions, and one-dimensional estimators of the mixture based on characteristic functions (see [33] and [37] for the second idea). We shall compare our random-projection method (RP) with the expectation-maximization (EM) algorithm.
There is a large literature on this subject. See for instance the well-known work in [21], the recent work [7] and the references therein. The article [21] considers general Gaussian multivariate models being $\epsilon$-statistically learnable, that is, $F=\sum_{i=1}^{n}\lambda_{i}F_{i}$ satisfying $\min_{i}\lambda_{i}\geq\epsilon$ and $\min_{i\neq j}\frac{1}{2}\int|f_{i}–f_{j}|\geq\epsilon$. In [7], very strong results are shown for the location case, where the covariances are all equal and known. In contrast, we just prove strong consistency, but this is done for the general case where the covariances are unknown and may be different, in an almost universal framework assuming only a differentiability condition on the characteristic functions, stated before Theorem 4.1 below.
4.2. Description of the algorithm
Let $P$ be a Gaussian mixture on $\mathbb{R}^{d}$, with density
$\sum_{j=1}^{n}{\lambda_{j}}\frac{1}{(2\pi\det(\Sigma_{j}))^{d/2}}\exp\Bigl{(}\frac{1}{2}(x-\mu_{j})^{T}\Sigma_{j}^{-1}(x-\mu_{j})\Bigr{)}\quad(x\in\mathbb{R}^{d}),$
where $\lambda_{1},\dots,\lambda_{n}\in[0,1]$ with $\sum_{j=1}^{n}\lambda_{j}=1$, where $\mu_{1},\dots,\mu_{n}\in\mathbb{R}^{d}$, and where $\Sigma_{1},\dots,\Sigma_{n}$ are $d\times d$ positive-definite matrices.
We write $\Lambda:=(\lambda_{1},\dots,\lambda_{n})$ and $\bm{\mu}:=(\mu_{1},\ldots,\mu_{n})$ and $\bm{\Sigma}:=(\Sigma_{1},\ldots,\Sigma_{n})$. Let $\Theta:=(\Lambda,\bm{\mu},\bm{\Sigma})$ denote the set of parameters.
Given a unit vector $u\in\mathbb{R}^{d}$ and $\theta\in\Theta$, we denote by $P_{\langle u\rangle}(\theta)$ the projection of the Gaussian mixture with parameters $\theta$ in the direction $u$. Given an iid sample of the Gaussian mixture $X_{1},\ldots,X_{N}$, let $P_{N,\langle u\rangle}$ be the projection of the empirical measure in the direction $u$.
Fix $k\geq(1/2)(\ell+m-1)(d^{2}+d-2)+1$, and let $u_{1},\ldots,u_{k}$ be iid random directions in $\mathbb{R}^{d}$.
We perform a two-step sequential estimation algorithm. In the first step, the vector $\Lambda$ is estimated. In the second step, the remaining parameters of the mixture are estimated.
Step 1: As is well known, if the data is projected using a linear transformation and a Gaussian mixture model is fitted on the original data, then the density on the projection subspace is again a mixture of Gaussian variables with covariance matrices and mean vectors obtained by a linear transformation of the parameters estimated in the original scale of the data.
For each random direction $u\in\mathbb{R}^{d}$, we can fit a univariate Gaussian mixture of the projected data as follows. We consider the model given by $\sum_{j=1}^{n}\lambda_{j}N_{u}(\upsilon_{j},\sigma^{2}_{j})$, where $\upsilon:=(\upsilon_{1},\dots,\upsilon_{n})$ and $\sigma^{2}:=(\sigma^{2}_{1},\dots,\sigma_{n}^{2})$ are the means and the variances of the projected distributions, given by $\upsilon_{j}:=\langle u,\mu_{j}\rangle,\ \sigma^{2}_{j}:=\langle u,\Sigma_{j}u\rangle\ (j=1,\dots n).$ Thus the distribution of the projected data is given by $\sum_{j=1}^{n}\lambda_{j}N\Bigl{(}\langle u,\mu_{j}\rangle,\langle u,\Sigma_{j}u\rangle\Bigr{)}$, and its characteristic function is
$\psi_{u}(t)=\psi_{u}(t,\upsilon,\sigma^{2},\Lambda):=\sum_{j=1}^{n}\lambda_{j}\exp[i\langle u,\mu_{j}\rangle t+\textstyle\frac{1}{2}\langle u,\Sigma_{j}u\rangle t^{2}].$
We denote by $\widehat{\psi}_{u}$ the empirical characteristic function (ECF) in the direction $u$, that is, $\widehat{\psi}_{u}(t)=\frac{1}{N}\sum_{j=1}^{N}\exp(i\langle u,X_{j}\rangle t)$. As in [33] and [37], we use the ECF procedure to estimate the parameters of mixtures of normal distributions. Roughly speaking, the objective is to minimize the distance between the characteristic function of the model and its empirical counterpart in the data. In the example, we consider a set of fixed points of the grid $t_{1},\dots,t_{m}$ where both functions are compared, see [33].
Let
$Z_{u}$ $:=\Bigl{(}\operatorname{Re}\psi_{u}(t_{1}),\ldots,\operatorname{Re}\psi_{u}(t_{m}),\operatorname{Im}\psi_{u}(t_{1}),\dots,\operatorname{Im}\psi_{u}(t_{m})\Bigr{)}^{t},$
$F_{u}(\upsilon,\sigma^{2},\Lambda)$ $:=\Bigl{(}\operatorname{Re}\widehat{\psi}_{u}(t_{1}),\dots,\operatorname{Re}\widehat{\psi}_{u}(t_{m}),\operatorname{Im}\widehat{\psi}_{u}(t_{1}),\dots,\operatorname{Im}\widehat{\psi}_{u}(t_{m})\Bigr{)}^{t},$
where $\operatorname{Re}z$ and $\operatorname{Im}z$ denote the real and imaginary parts of the complex number $z$ respectively. We define the estimator $(\widehat{\upsilon}_{u,(1)},\widehat{\sigma}^{2}_{u,(1)},\widehat{\Lambda}_{u,(1)})$ as the values of the parameters that minimize the product
$\bigl{(}Z_{u}-F_{u}(\upsilon,\sigma^{2})\bigr{)}^{t}W\bigl{(}Z_{u}-F_{u}(\upsilon,\sigma^{2})\bigr{)},$
where $W$ is some weighting matrix defined so as to ensure that the sample orthogonality conditions are made as close as possible to zero. The optimal matrix $W$ can be estimated by means of a heteroskedasticity and consistent autocorrelation (HAC) matrix such as the one proposed in [22]. This is just the generalized moment estimator, see [15], where the moment conditions are of the form $h(t,u,X_{j},\upsilon,\sigma^{2})=\exp(it\langle u,X_{j}\rangle)-\psi_{u}(t,\upsilon,\sigma^{2})$, and obviously $E(h(t_{i},u,X_{j},\upsilon,\sigma^{2},\Lambda))=0$ for all $i=1,\dots,m.$
The procedure is replicated in $k$ random directions $u_{1},\dots,u_{k}$, and the estimator of $\Lambda$ is $\widehat{\Lambda}_{(1)}=\frac{1}{k}\sum_{r=1}^{k}\widehat{\Lambda}_{u_{r},(1)}$.
Step 2: Replacing $\Lambda$ by its estimated value $\widehat{\Lambda}_{(1)}$, we repeat Step 1, but only for the parameters $(\upsilon_{u},\sigma^{2}_{u})$. The estimators
$\widehat{\upsilon}_{u_{r},(2)}$ $:=(\widehat{\upsilon}_{u_{r},(2),1},\dots,\widehat{\upsilon}_{u_{r},(2),n}),\,\,\widehat{\sigma}^{2}_{u_{r},(2)}:=(\widehat{\sigma}^{2}_{u_{r},(2),1},\dots,\widehat{\sigma}^{2}_{u_{r},(2),n}),$
are obtained for $r=1,\dots,k$. Then the estimators for the mixture parameters can be obtained by least squares:
(3) $\widehat{\mu}_{j}:=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\Bigl{(}\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},(2),j}\Bigr{)}^{2},$
(4) $\widehat{\Sigma}_{j}:=\operatorname*{argmin}_{\Sigma_{j}}\sum_{r=1}^{k}\Bigl{(}\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\Bigr{)}^{2},$
for all $j=1,\dots,n$.
Remark.
(i) If $\Lambda$ is known or estimated by another mechanism, then the estimation procedure starts at Step 2. For example, as in [5], this happens when one starts with a cluster procedure that has already been performed. This allows to have a $\Lambda$-estimator across the group sizes obtained.
(ii) Robust estimators can be obtained by replacing the $L^{2}$-minimization by the $L^{1}$-one, that is, by solving the equations
$\widehat{\mu}_{j}=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\bigl{|}\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},j}\bigr{|},\ \ \widehat{\Sigma}_{j}=\operatorname*{argmin}_{\Sigma_{j}}\sum_{r=1}^{k}\bigl{|}\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},j}\bigr{|},$
for all $j=1,2,\dots,n$.
Alternatively, in the spirit of the proposed median of means, see for instance [6] and the references therein, one can proceed as as follows. We split the sample into $L$ disjoint subsamples. For each subsample and $j$ we calculate $\widehat{\mu}_{j,\ell}$ and $\widehat{\Sigma}_{j,\ell}$, $j=1,\dots n$, $\ell=1,\ldots,L$ according to equations (3) and (4), and we take the median of the corresponding $L$ values obtained for each $j$. Then we proceed as before.
(iii) We have to make sure that the estimators $\Sigma_{j}$ are positive semi-definite matrices, by using semidefinite optimization as follows (see for instance [17]).
Let $\mathcal{S}^{d}$ denote the space of real symmetric $d\times d$ matrices, equipped with the usual inner product $\langle S_{1},S_{2}\rangle:=\operatorname{trace}(S_{1}^{t}S_{2})$, and let $\mathcal{S}^{+,d}$ be the closed convex cone consisting of positive semi-definite matrices. Let $\Pi_{\mathcal{S}^{+,d}}$ be the projection of $\mathcal{S}^{d}$ onto $\mathcal{S}^{+,d}$, given by $\Pi_{\mathcal{S}^{+,d}}(S):=\operatorname*{argmin}_{S^{\prime}\in\mathcal{S}^{+,d}}\|S^{\prime}-S\|\quad(S\in\mathcal{S}^{d})$. It turns out that, if the spectral decomposition of $S\in\mathcal{S}^{d}$ is given by $S=U\Gamma U^{t}$, where $\Gamma=\operatorname{diag}(\gamma_{1},\dots,\gamma_{d})$, then $\Pi_{\mathcal{S}^{+,d}}(S)=U\Gamma^{+}U^{t}$, where $\Gamma^{+}=\operatorname{diag}(\gamma_{1}^{+},\dots,\gamma_{d}^{+})$ and $\gamma_{j}^{+}=\max(\gamma_{j},0)\leavevmode\nobreak\ (j=1,\dots,d)$.
If we want a strictly positive-definite matrix, then we can add a term $\delta I_{d}$, where $I_{d}$ is the $d\times d$ identity matrix and $\delta$ is a very small positive constant.
(iv) Lastly, we note that the algorithm can also be applied to mixtures of multivariate $t$-distributions.
4.3. A consistency result
The results in [33] and [37] that we will use below are based on a general theorem in [16], where the only assumption is that the characteristic function is differentiable. More precisely, it is assumed that the regular case holds, namely that $I_{n}(t)$ can be differentiated under the integral sign.
Theorem 4.1.
Under the weak assumption that,
[…]
the estimators $\widehat{\Lambda}_{(1)},\widehat{\mu}_{j},\widehat{\Sigma}_{j}$ for $j=1,\dots,n$ are strongly consistent.
Proof.
First observe that, from the results in [33] and [37] for the finite set of estimators for the projected one-dimensional projections, we can derive the strong consistency of $\widehat{\Lambda}_{(1)}$. See [33, Proposition 1], where strong consistency and asymptotic normality are established.
For the next step, using the value of $\widehat{\Lambda}_{(1)}$, we consider the estimators of $\widehat{\mu}_{j}$ and $\widehat{\Sigma}_{j}$ for $j=1,\dots n$ given by the least-squares equations (3) and (4). Since
$\sum_{r=1}^{k}\bigl{(}\langle u_{r},\mu_{j}\rangle-\upsilon_{u_{r},j}\bigr{)}^{2}=0\quad\text{and}\quad\sum_{r=1}^{k}\bigl{(}\langle u_{r},\Sigma_{j}u_{r}\rangle-\sigma^{2}_{u_{r},j}\bigr{)}^{2}=0$
for the true values of $\mu_{j}$ and $\Sigma_{j}\leavevmode\nobreak\ (j=1,\dots n)$, we have
[…]
and
[…]
for all $j=1,\dots,n$. This concludes the proof. ∎
4.4. Simulations
4.4.1. Example 1.
Using the above algorithm, we estimate the parameters of a mixture of two bivariate t-distributions $\textrm{t}\left(\eta,\mu,\Sigma\right)$ with $\eta$, $\mu$ and $\Sigma$ the degrees of freedom, location and scale matrix parameters respectively. Consider
$F:=\lambda_{1}\textrm{t}(\eta,\mu_{1},\Sigma_{1})+(1-\lambda_{1})\textrm{t}(\eta,\mu_{2},\Sigma_{2}),$
with
$\mu_{1}=(1,0),\quad\mu_{2}=(5,0),\quad\Sigma_{1}=\begin{pmatrix}1&0\\ 0&1/2\end{pmatrix},\quad\Sigma_{2}=\begin{pmatrix}1/2&0\\ 0&1\end{pmatrix},\quad\lambda_{1}=0.3\textrm{ and }\eta=4.$
From $1000$ iid simulations of $F$, see Figure 5, the objective is to estimate the parameters $\mu_{1}$, $\mu_{2}$, $\Sigma_{1}$ and $\Sigma_{2}$. We implement the proposed algorithm by projecting the data in $k=50$ directions, chosen at random. This procedure is replicated $100$ times.
Figure 5. Depiction of $1000$ simulations of the mixture $F$.
In this Example, we present a small comparative study of the estimation derived by the random projections method (RP) with respect to the expectation-maximization method for mixtures of t-distributions (we denote EM-st). This last method, developed in [24], is implemented by means of the function fmmst of the EMMIXuskew package of the R language. In this case the skew parameter was prefixed to 00.
Table 1 shows the average estimates of the parameters in the replicates with their respective deviations. Here, we can clearly observe the adequate fitting of our estimators (at least within this brief example).
Table 1. Averages values of the estimated parameters over $100$ replicates of the algorithm. Marginal standard deviations (sd) are shown in parentheses.
Algorithm Component $\mu_{1}$ $\mu_{2}$ $\sigma_{11}$ $\sigma_{22}$ $\sigma_{21}$
Average estimated values
Values 1 1 0 1 1/2 0
2 5 0 1/2 1 0
RP 1 0.99 -0.01 0.99 0.42 -0.01
sd (0.10) (0.10) (0.16) (0.15) (0.08)
2 4.99 0.014 0.55 0.67 -0.01
sd (0.13) (0.12) (0.06) (0.21) (0.08)
EM-st 1 1.22 0.07 0.89 0.62 -0.01
sd (0.15) (0.16) (0.14) (0.13) (0.09)
2 4.56 -0.10 0.51 1.15 -0.01
sd (0.10) (0.25) (0.07) (0.15) (0.07)
In each replicate, we can calculate the $L^{2}$-distance (error) between the estimated mean and mean vectors and their respective true values, see Figure 6. The estimation of the mixture parameter is similar, giving on average $\widehat{\lambda_{1}}_{EM-st}=0.312$ and $\widehat{\lambda_{1}}_{RP}=0.293$ for EM-st and RP respectively. The code shows the convergence of the EM algorithm and similar execution times in both algorithms.
Figure 6. Comparison between the EM-st and RP methods using the $L^{2}$-distance between the parameters.
4.4.2. Example 2
We repeat the previous example, but now simulating the data for the mixture of two Gaussian distributions, contaminated with a noise from a bivariate Cauchy distribution centered at $\mu_{3}$ (which we denote by $\mathcal{C}a\left(\mu_{3}\right)$). Thus
$F=\lambda_{1}N(\mu_{1},\Sigma_{1})+(1-\lambda_{1}-\gamma)N(\mu_{2},\Sigma_{2})+\gamma\mathcal{C}a\left(\mu_{3}\right),$
with
$\mu_{1}=(1,0),\quad\mu_{2}=(5,0),\quad\mu_{3}=(3,0),\quad\Sigma_{1}=\begin{pmatrix}4&0\\ 0&2\end{pmatrix},\quad\Sigma_{2}=\begin{pmatrix}2&0\\ 0&4\end{pmatrix},\quad\lambda_{1}=0.3\textrm{ and }\gamma=0.05.$
Here $\gamma$ indicates the proportion of outliers. In this case we re-estimate the original mixture parameters, and the $L^{2}$-errors are plotted in Figure 7.
In this framework, our method is compared to a mixture estimation using a robust variant of EM (which we denote by RobEM). These estimators are introduced in [13] and are implemented in the RGMM package of the R language. The methodology they propose consists of modifying the EM algorithm, in particular the M-step, where they replace the mean and variance estimates by robust versions derived from the median and the median of the covariance matrix, respectively.
Figure 7 shows that, in this example, our method presents in general smaller errors in the estimation of the parameters than RobEM.
Figure 7. Comparison between the Robust EM and RP Boxplots when outliers are added, taking $\mu_{3}=(3,0)$.
4.5. An example with real data
The relationship between school performance and students’ socioeconomic and cultural status has been widely studied, see [36]. This association is even stronger in the Latin American countries, see [8].
In Uruguay, this trend can be seen in the educational evaluation studies of the National Institute for Educational Evaluation (INEEd). A representative sample of $6437$ students in the third year of secondary education in the country in 2022 is considered. The sample is comprised of students attending two types of educational institutions: Public (free and funded by the state) and Private (depending on the payment of tuition by their students or private sources of financing). In the sample, $4852$ and $1585$ students attend public and private schools, respectively. By means of a multiple-choice test, using Item Response Theory, a score in Mathematics is assigned to the item. In addition, an index of each student’s socioeconomic and cultural level is constructed from data collected in a personal questionnaire. [^2] The database and indexes mentioned above are open and available at https://www.ineed.edu.uy/nuestro-trabajo/aristas/
Figure 8 depicts a bivariate histogram in hexagonal cells. This Figure shows the positive relationship between the two measures, where an asymmetric distribution is observed.
Figure 8. Hexagonal binning for bivariate data. The plane is tessellated by a regular grid of hexagons (cells). Then the hexagons are plotted using an intensity of color proportional to the number of points falling in each cell. Higher socio-economic level for smaller values on the $x$-axis, better mathematical ability for larger values on the $y$-axis.
The objective is to model this bivariate distribution by means of a mixture of normals and to explain a possible reason for this asymmetry. Using the method that we have developed, the following estimated parameters of a mixture of two Gaussians are obtained: $\widehat{\lambda}_{1}=0.75,\ \widehat{\mu}_{1}=(-0.14,-0.15),\ \widehat{\mu}_{2}=(1.16,0.47)$, and
$\widehat{\Sigma}_{1}=\begin{pmatrix}0.28&0.05\\ 0.05&0.45\end{pmatrix},\quad\widehat{\Sigma}_{2}=\begin{pmatrix}0.16&0.07\\ 0.07&0.40\end{pmatrix}.$
These two data patterns are possibly explained by the type of institution that the students attend. In this example, we consider $\lambda$ to be known a priori, i.e., the proportion of students attending Public educational institution is $0.75$. If we calculate the vector of means and the matrix of empirical variances and covariances of the sample conditioned to each institution, we obtain $\widehat{\mu}_{\text{Public}}=(-0.16,-0.180),\ \widehat{\mu}_{\text{Private}}=(1.18,0.51)$, and
$\widehat{\Sigma}_{\text{Public}}=\begin{pmatrix}0.45&0.10\\ 0.10&0.75\end{pmatrix},\quad\widehat{\Sigma}_{\text{Private}}=\begin{pmatrix}0.20&0.12\\ 0.12&0.59\end{pmatrix}.$
These values are similar to those obtained in the mixture estimation.
5. Conclusions and future work
We consider two important statistical problems: that of testing and measuring the agreement between two different random partitions for cluster analysis, and that of estimating for mixtures of multivariate normal distributions and mixtures of $t$-distributions based of univariate projections.
A model-based measure of agreement is introduced and compared with more classical measures. Instead of looking at the number of matchings, we look at the distance between the mixture distributions associated with them, which provide much more information about the agreement between the partitions.
We propose a new estimator for multivariate mixture distributions based on one-dimensional projections and their respective characteristic functions. We derive the strong consistency of them and in a small simulation study we compare with the more classical EM-type estimators, where we observe a very good behaviour.
The relationship between school performance and students’ socioeconomic and cultural status is analyzed in a real data example. We find two different patterns that are possible, explained by the type of institution that the students attend : public or private.
Rates of convergence and the asymptotic distribution of the estimators remain as open problems that are not addressed in the present manuscript.
The number of directions $k$ into which we project is not a smoothing parameter that should be chosen in an optimal way. In practice we suggest to use a larger value for $k$ (without increasing too much the computational time) than the one provided by Theorem 2.1 which corresponds to the case when the mixtures are known, whereas we are dealing with empirical distributions which are close to the unknown underlying distributions but not exactly equal. Depending on the problem, among the directions chosen at random it may appear some directions which are close to bad ones. Increasing the value of $k$ solves the problem. In our experience after a reasonable value the performance stabilizes. We illustrate this fact in Figure 3 .
Also some more computationally oriented work, which includes different mixtures of distributions, sample sizes and number of clusters, is also not included in the present manuscript.
Acknowledgments
Fraiman and Moreno were supported by grant FCE-3-2022-1-172289 from ANII (Uruguay), Program for the Development of Basic Sciences-PEDEClBA (Uruguay) and 22520220100031UD from CSIC (Uruguay). Ransford was supported by NSERC Discovery Grant RGPIN-2020-04263.