### v4

Abstract
We address two important statistical problems: that of estimating for mixtures of multivariate normal distributions and mixtures of $t$-distributions based on univariate projections, and that of measuring the agreement between two different random partitions. The results are based on an earlier work of the authors, where it was shown that mixtures of multivariate Gaussian or $t$-distributions can be distinguished by projecting them onto a certain predetermined finite set of lines, the number of lines depending only on the total number of distributions involved and on the ambient dimension. We also compare our proposal with robust versions of the expectation-maximization method EM. In each case, we present algorithms for effecting the task, and compare them with existing methods by carrying out some simulations.
keywords:
clustering, Gaussian distribution, identifiability, mixture, projection, random partition, $t$-distribution
pacs:
pacs:
JEL Classification]D8, H51
MSC Classification]35A01, 65L10, 65L12, 65L20, 65L70
1 Introduction
1.1 Random projections in statistics
Random-projection (RP) techniques have become a cornerstone in high-dimensional statistical analysis. Over the last two decades, the theoretical and applied developments of RP have flourished. Early applications emphasized computational efficiency and scalability, particularly for clustering [8], matrix approximations, [70, 41], and kernel learning [65]. The work presented in [1] introduced more efficient sparse and structured random matrices, which proved crucial for large-scale implementations.
The best-known method for applications involving random projections is the one based on the Johnson–Lindenstrauss lemma, which has numerous applications, especially in image processing. Although there are many other proposals based on RP, in what follows we will briefly describe the Johnson–Lindenstrauss approach below and then focus solely on methods based on extensions of the Cramér–Wold theorem. We will describe each of them in more detail below.
1.1.1 RP via the Johnson–Lindenstrauss lemma
This lemma can be stated as follows.
Lemma (Johnson–Lindenstrauss lemma, [46]). Let $\aleph_{n}$ be a set of $n$ points in $\mathbb{R}^{d}$, let $0<\epsilon<1$ and let $q>(8\log n)/\epsilon^{2}$. Then there exists a linear map $f:\mathbb{R}^{d}\to\mathbb{R}^{q}$ such that, for all $u,v\in\aleph_{n}$,
$(1-\epsilon)\|u-v\|^{2}\leq\|f(u)-f(v)\|^{2}\leq(1+\epsilon)\|u-v\|^{2}.$ (1)
Thus any set of points in high-dimensional Euclidean space can be linearly embedded into a much lower-dimensional space such that pairwise distances are approximately preserved. This result lays the theoretical foundation for the widespread application of RP in modern data science and machine learning.
A natural question that arises is how to find the linear map $f$. Dasgupta and Gupta [23] propose a solution based on the following result, which allows us to find $f$ in randomized polynomial time.
Theorem 1 (Dasgupta–Gupta [23]).
Let $\aleph_{n}$ and $\epsilon$ be as in Theorem 1, and let $q>(4\log n)/(\epsilon^{2}/2-\epsilon^{3}/3)$. Define $f(u):=\sqrt{(d/q)}PUu$, where $U$ is a $d\times d$ unitary matrix whose columns are iid (independent and identically distributed) random vectors uniformly distributed on the unit sphere in $\mathbb{R}^{d}$, and $P:\mathbb{R}^{d}\to\mathbb{R}^{q}$ denotes the projection onto the first $q$ coordinates. Then
$\mathbb{P}\Bigl((1-\epsilon)\|u-v\|^{2}\leq\|f(u)-f(v)\|^{2}\leq(1+\epsilon)\|u-v\|^{2})~\forall u,v\in\aleph_{n}\Bigr)\geq 1/n.$
Recent studies continue to build on the Johnson–Lindenstrauss lemma, investigating its application in modern contexts, for example, as a method for dimensionality reduction [22], particularly in machine-learning applications [30], in high-dimensional clustering [2], and nearest neighbor search [49]. These studies demonstrate how RP can preserve important properties in high-dimensional spaces while dramatically reducing the computational complexity of various algorithms.
A nice property of the inequality (1) is that it does not depend on the dimension $d$. However, if $\epsilon$ is small then $q$ becomes very large.
1.1.2 RP via the Cramér–Wold theorem
A more recent and statistically grounded direction emerged in the work of Cuesta, Fraiman and Ransford [18], based on extensions of the Cramér–Wold theorem.
In its original form, established in [15], the Cramér–Wold theorem says that a multivariate distribution is uniquely determined by its one-dimensional projections. Subsequent work of Rényi [67], Gilbert [38], Bélisle–Massé–Ransford [6] and Cuesta–Fraiman–Ransford [18] showed that, under appropriate assumptions on the growth of its moments, a multivariate distribution is determined by a ‘sufficiently rich’ set of lower-dimensional projections, thereby allowing it to be determined from random projections onto low-dimensional subspaces.
Here is one such result. In what follows, given a Borel probability measure $P$ on $\mathbb{R}^{d}$ and a vector subspace $H$ of $\mathbb{R}^{d}$, we denote by $P_{H}$ the projection of $P$ onto $H$, defined as the Borel probability measure on $H$ given by
$P_{H}(B):=P(\pi_{H}^{-1}(B)),$
where $\pi_{H}:\mathbb{R}^{d}\to H$ is the orthogonal projection of $\mathbb{R}^{d}$ onto $H$. Also, if $x\in\mathbb{R}^{d}$, then $\langle x\rangle$ denotes the vector subspace spanned by $x$.
Theorem 2 (Cuesta–Fraiman–Ransford [18, Corollary 3.2]).
Let $P$ and $Q$ be Borel probability measures on $\mathbb{R}^{d}$, where $d\geq 2$. Assume that:
(i) the absolute moments $m_{n}=\int\|x\|^{n}\,dP(x)$ are finite and satisfy the condition $\sum_{n=1}^{\infty}m_{n}^{-1/n}=\infty$;
(ii) the set $\{x\in\mathbb{R}^{d}:P_{\langle x\rangle}=Q_{\langle x\rangle}\}$ is of positive Lebesgue measure.
Then $P=Q.$
Cuesta–Fraiman–Ransford focused on using RP not only as a computational tool but as a vehicle for inferring robust statistical structure. In particular, they demonstrated how statistical procedures random-projection tests can preserve desirable properties such as consistency and robustness while remaining computationally feasible.
This line of work differs fundamentally from algorithmic RP approaches. Rather than using RP solely for dimensionality reduction, Cuesta, Fraiman and their collaborators exploit the probabilistic distribution of projections to characterize centrality, perform goodness-of-fit tests [16], depth measure [20], detect outliers [56], and hypothesis tests [36]. For example, in their framework, distributions are projected onto randomly chosen lines, and classical univariate techniques are applied. Repeating this process over multiple projections yields statistical conclusions that are consistent with high-dimensional structures.
These methods have inspired extensions in several applied domains. For instance, random-projection depth was later adapted to functional data analysis, see [21] and [19]. The article [60] proposes tests for functional data using projections, while [4] explores robust estimators via projection-pursuit methods. These studies reinforce the versatility of random projections as not only an algorithmic trick but a robust statistical principle. Moreover, this framework aligns conceptually with the work of [26], which emphasized the geometric underpinnings of high-dimensional phenomena, although the focus in [26] is often more combinatorial.
In contrast to mainstream applications of random projection, such as kernel approximation [65] or compressed sensing [11], the approach developed in [17, 18] is distinguished by its nonparametric nature and inferential focus. Rather than focusing on data transformations that preserve structural relationships for machine learning, the latter work is more oriented to preserving the geometric properties of the data for inferential purposes. This distinction in focus underscores the role of RP in nonparametric inference, as it provides an intuitive tool for deriving statistical conclusions from data without imposing stringent parametric assumptions.
More recently, there have been some further refinements of the Cramér–Wold theorem showing that, in some cases, just finitely many one-dimensional projections suffice to determine the measure. For example, the following result shows that, in the case of elliptical distributions (which includes Gaussian and $t$-distributions), just $(d^{2}+d)/2$ projections suffice.
Theorem 3 (Fraiman–Moreno–Ransford [35]).
Let $v_{1},\dots,v_{d}$ be linearly independent vectors in $\mathbb{R}^{d}$, and let $S=\{v_{j}+v_{k}:1\leq j\leq k\leq d\}$. If $P,Q$ are elliptical distributions in $\mathbb{R}^{d}$ and if $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
In [34] these results were extended to the case of certain multivariate mixtures. It was shown that mixtures of multivariate Gaussian or $t$-distributions can be distinguished by projecting them onto a certain predetermined finite set of lines, the number of lines depending only on the total number of distributions involved and on the ambient dimension. This last result is described in more detail in §2, and its applications form the subject of the rest of the paper.
1.1.3 On the pros and cons of RP-methods
Despite its many advantages, the RP methodology is not without its limitations. One notable drawback is that while it offers dimensionality reduction, it does so at the expense of certain properties of the data. Although the Johnson–Lindenstrauss lemma guarantees approximate distance preservation, it does not ensure exact preservation, especially when data points lie on complex, nonlinear manifolds. This means that RP may distort some aspects of the data distribution, especially for applications like clustering or classification, where fine-grained distinctions between data points are critical, see [8].
Moreover, while RP has been shown to be robust in certain high-dimensional settings, its performance can be sensitive to the type of distribution or structure inherent in the data. For example, in datasets with heavy tails or outliers, RP techniques may not preserve the underlying distribution as effectively as more robust methods like those based on depth functions, see [20]. Thus, while RP is often computationally efficient, it may sometimes fail to capture subtle patterns that are crucial in some statistical problems.
A significant challenge also arises from the inherent randomness of RP. This randomness introduces variability in the performance of algorithms across different projections. While multiple projections typically improve the consistency of results, determining the number of projections necessary for a reliable outcome is still an open question in many applications, see [18]. This issue has spurred research on the optimal number of projections required for specific tasks, particularly in the context of robust statistics and outlier detection.
Despite these challenges, the advantages of RP remain substantial. Its ability to reduce dimensionality while retaining much of the essential structure of the data makes it a powerful tool in large-scale statistical learning. Moreover, the simplicity of random projections allows for faster computations compared to other dimensionality reduction techniques, such as principal component analysis (PCA) or t-SNE, which can be computationally expensive in high-dimensional spaces, see [78].
As modern datasets continue to grow in complexity and dimensionality, the statistical insight provided by projection-based depth and testing becomes increasingly relevant. It offers a complement—and at times a counterpoint—to the more engineering-oriented uses of RP found in neural networks [14], sketching algorithms [50], or NLP embeddings [77]. The blend of theoretical rigor and practical robustness in the works of Cuesta, Fraiman, Ransford and their collaborators provides tools that are especially suitable for applied statisticians working under real-world constraints, such as noisy data or nonstandard distributions.
Although the manuscript addresses (i) parameter estimation for multivariate mixtures and (ii) agreement between random partitions, both contributions are governed by the same principle: within the Gaussian/$t$-mixture classes considered here, the underlying distribution is determined by a finite collection of one-dimensional projections along a strong sm-uniqueness set (Section 2). Section 3 exploits this fact constructively for estimation, whereas Section 4 exploits it comparatively to quantify agreement by aggregating projected discrepancies.
Notation.
Throughout the manuscript, $m$ denotes the number of mixture components, $d$ the ambient dimension, $N$ the sample size (i.i.d. observations), and $k$ the number of projection directions used by the random-projection procedures. In particular, we use $N$ for the sample size to avoid notational conflicts with other indices.
1.2 Roadmap of the rest of the article
In this article we consider two problems: that of estimating for mixtures of multivariate normal distributions and $t$-distributions based on univariate projections, and that of measuring the agreement between two random partitions (the output of two different model–based clustering procedures). To solve these problems we make use of some of the results given in [34].
Section 2 begins with a short introduction to the notion of the Cramér–Wold device, and the main notions and results to be used are described.
In Section 3, we present a new procedure for estimating mixtures of multivariate normal distributions, and mixtures of $t$-distributions based on one-dimensional projections of the mixture based on characteristic functions (ECF), see [76] and [82], and using those univariate mixtures to provide consistent estimators of the parameters of the multivariate mixture.
The proposed random-projection method (RP) is compared with the well-known expectation-maximization method (EM) and robust versions of EM in some simulations. We use some results in [76] (see also [85] and [82]) for some asymptotic properties. An important fact, mentioned in [76], is that “It is shown that, using Monte Carlo simulation, the finite sample properties of the ECF estimator are very good, even in the case where the popular maximum likelihood estimator fails to exist…”. We conclude with a further example using real data from the National Institute for Education Evaluation (INEEd) in Uruguay. The dataset is open access, and the codes, written in R, are available on GitHub [^1] https://github.com/mrleomr/MultivariateMixture or can be requested directly from the authors.
In Section 4 we discuss the problem of comparing random partitions by model-based methods. Comparing partitions is an important problem in particular in cluster analysis, and a large amount of work on the subject introduced several different proposals. The most popular measures are given by the Rand index [66], the adjusted Rand index [45], the Jaccard index, the Dunn index, the silhouette index, and the Xie–Beni index among others. See for instance the reviews by Arabie and Boorman [3], Fowlkes and Mallows [33], among others.
The way that we approach this problem is to measure the agreement between the probability measures that have produced those partitions (the output of two different model-based clustering procedures). Model-based approach provides a good interpretation of this measure in terms of the underlying unknown distributions. Indeed, if the mixtures distributions are close, then we can conclude that the partitions are close, even when they have been produced by different data bases, providing a more general setup to the problem.
Algorithms implementing these ideas are described, and then illustrated with some simulations. The simulations indicate a high degree of agreement between several well-known distance-measures and the one that we propose.
Our arguments rely on the notion of a strong symmetric-matrix uniqueness set (strong sm-uniqueness set), which provides a finite Cramér–Wold system for Gaussian and $t$-mixtures. We recall the definition and its construction in Section 2; see [35, 34].
2 Some preliminaries and recent results to be used
In this section, we introduce the fundamental concepts necessary for understanding our work. In particular, we clarify the notion of mixture, and describe a version of the Cramér–Wold theorem for mixtures of Gaussian or Student distributions.
2.1 Mixtures
Mixture models are fundamental tools in statistics and machine learning, providing a flexible framework for representing heterogeneous data as combinations of simpler component distributions. Since the pioneering work of Pearson [61], they have been extensively used in tasks such as clustering [74], density estimation [48], and the modeling of latent structures [59]. Finite mixture models have also found applications across a broad range of disciplines, including astronomy, biology, genetics, economics, social sciences, and engineering [52].
Estimation rates for finite mixture distributions, that is, for the mixture parameters, are also studied in [12, 44]. A more detailed review of the literature, including specific cases involving Gaussian or Student $t$ mixture distributions, can be found in [51].
Let $d\geq 1$ and let $\mathcal{P}$ be a family of Borel probability measures on $\mathbb{R}^{d}$. A $\mathcal{P}$-mixture is defined as a convex combination of measures from $\mathcal{P}$. In other words, it takes the form
$\sum_{j=1}^{n}\lambda_{j}P_{j}$
where $P_{1},\dots,P_{n}\in\mathcal{P}$ and $\lambda_{1},\dots,\lambda_{n}\geq 0$ with $\sum_{j=1}^{n}\lambda_{j}=1$.
Once the data has been projected, we will focus on univariate distributions. Specifically, if $d=1$, we will consider the two families of univariate distributions: Gaussian distributions and $t$-distributions. Gaussian distributions are characterized by their densities
$f_{\mu,\sigma}(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\Bigl(-\frac{(x-\mu)^{2}}{2\sigma^{2}}\Bigr)\quad(x\in\mathbb{R}).$ (2)
A $t$-distribution on $\mathbb{R}$ with $\nu$ degrees of freedom is a Borel measure characterized by the density function
$f_{\nu,\mu,\sigma}(x)=c_{\nu,\mu,\sigma}\left(1+\frac{(x-\mu)^{2}}{\nu\sigma^{2}}\right)^{-(\nu+1)/2},$
where $\nu$ is a positive integer, $\mu\in\mathbb{R}$ and $\sigma>0$. The constant $c_{\nu,\mu,\sigma}$ is chosen to ensure that
$\int_{\mathbb{R}}f_{\nu,\mu,\sigma}(x)\,dx=1.$
This distribution has a mean of $\mu$ (provided that $\nu>1$) and a variance of $\sigma^{2}\nu/(\nu-2)$ (if $\nu>2$).
In the multivariate case, where $d>1$, a Gaussian measure $P$ on $\mathbb{R}^{d}$ is defined by a density of the form
$\frac{1}{(2\pi\det(\Sigma))^{1/2}}\exp\Bigl(-\frac{1}{2}(x-\mu)^{T}\Sigma^{-1}(x-\mu)\Bigr)\quad(x\in\mathbb{R}^{d}),$ (3)
where $\mu\in\mathbb{R}^{d}$, and where $\Sigma$ is a real $d\times d$ positive-definite matrix.
A Gaussian mixture is a measure on $\mathbb{R}^{d}$ that represents a finite convex combination of Gaussian measures. Mixtures of multivariate Gaussian distributions exhibit several advantageous properties. In particular, Titterington et al. [75] demonstrate that Gaussian kernel density estimators can approximate any continuous density given a sufficient number of kernels, establishing their universal consistency (see also Scott [72]). Additionally, it is well known that Gaussian mixtures are weak*-dense in the space of all Borel probability measures on $\mathbb{R}^{d}$.
Gaussian-mixture models have proven to be highly effective in modeling various real-world data. For a comprehensive examination of their properties, we refer to Titterington et al. [75]. These models are flexible and general, with relevant applications documented in numerous fields, including density estimation, machine learning, and clustering. As noted in [13], significant applications include Pearson’s work on modeling crab data [61], Roeder’s studies [69], and the research by Schork, Allison and Thiel in genetics [71], as well as Everitt’s investigations involving schizophrenia patients [29], among many others.
Estimating these models can be quite complex, particularly for high-dimensional data, and typically involves Markov-chain Monte-Carlo methods within a Bayesian framework. Important applications for cluster analysis are discussed in Tadesse, Sha, and Vannucci [73], as well as in Raftery and Dean [64].
Similarly, when the data exhibit heavy tails, a similar development is plausible using mixtures of Student distributions. A Student distribution ($t$-distribution) on $\mathbb{R}^{d}$ is a measure with density of the form
$f_{\nu,\mu,\Sigma}(x)=c_{\nu,\mu,\Sigma}\Bigl(1+\frac{(x-\mu)^{T}\Sigma^{-1}(x-\mu)}{\nu}\Bigr)^{-(\nu+d)/2}\quad(x\in\mathbb{R}^{d}),$
where $\nu$ is a positive integer, $\mu$ is a vector in $\mathbb{R}^{d}$, and where $\Sigma$ is a positive-definite $d\times d$ matrix. Once again, the constant $c_{\nu,\mu,\Sigma}$ is chosen to ensure that $\int_{\mathbb{R}^{d}}f_{\nu,\mu,\Sigma}(x)\,dx=1$.
2.2 A Cramér–Wold Theorem for Gaussian mixtures and $t$-mixtures
In this section, we address the problem for equality between two Gaussian mixtures by analyzing a finite number of projections. The basic Theorem supporting this approach consists of two key components. The first is the abstract result presented in Theorem 4.1 in [34]. The second is a characterization of Cramér–Wold systems for Gaussian measures in $\mathbb{R}^{d}$ (and, more generally, for elliptical distributions), established in [35, Theorems 1 and 2], which we will now summarize.
Let $S$ be a set of vectors in $\mathbb{R}^{d}$. The corresponding set of lines $\{\langle x\rangle:x\in S\}$ forms a Cramér–Wold system for the Gaussian measures in $\mathbb{R}^{d}$ if and only if $S$ satisfies the property that the only real symmetric $d\times d$ matrix $A$ for which $x^{T}Ax=0$ for all $x\in S$ is the zero matrix. A set $S$ with this property is referred to as a symmetric-matrix uniqueness set (or sm-uniqueness set for short).
In [35], it was demonstrated that an sm-uniqueness set for $\mathbb{R}^{d}$ must span $\mathbb{R}^{d}$and contain at least $(d^{2}+d)/2$ vectors. We will denote $S$ as a strong sm-uniqueness set if every subset of $S$ containing $(d^{2}+d)/2$ vectors is also an sm-uniqueness set.
We now recall the Cramér–Wold Theorem for Gaussian mixtures [34, Theorem 4.1].
Theorem 4.
Let $P$ and $Q$ be convex combinations of $m$ Gaussian measures on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(2m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
We also recall an analogue of Theorem 4 for mixtures of multivariate $t$-distributions, given in [34, Theorem 4.2], thereby allowing heavy-tailed distributions.
Theorem 5.
Let $P$ and $Q$ be convex combinations of $m$ multivariate $t$-distributions on $\mathbb{R}^{d}$ respectively. Let $S$ be a strong sm-uniqueness set for $\mathbb{R}^{d}$ containing at least $(1/2)(2m-1)(d^{2}+d-2)+1$ vectors. If $P_{\langle x\rangle}=Q_{\langle x\rangle}$ for all $x\in S$, then $P=Q$.
2.3 Strong sm-uniqueness sets
Theorems 4 and 5 beg the question as to whether there exist strong sm-uniqueness sets of arbitrarily large cardinality. An affirmative answer is given in the following result based on [34, Theorem 4.3], which also provides a realistic method for generating them.
We next recall the notion of a (strong) sm-uniqueness set, which is needed for the theorem below.
Definition 1 (sm-uniqueness and strong sm-uniqueness sets [35, 34]).
Let $S\subset\mathbb{R}^{d}$. We say that $S$ is an sm-uniqueness set if the only real symmetric $d\times d$ matrix $A$ satisfying $x^{\top}Ax=0$ for all $x\in S$ is $A=0$. If $|S|\geq d(d+1)/2$, we say that $S$ is strong if every subset of $S$ with cardinality $d(d+1)/2$ is itself an sm-uniqueness set.
Theorem 6.
Let $d\geq 2$, let $k\geq(d^{2}+d)/2$, and let $v_{1},\dots,v_{k}$ be independent random vectors in $\mathbb{R}^{d}$ whose distributions are given by densities on $\mathbb{R}^{d}$. Then, with probability one, the set $\{v_{1},\dots,v_{k}\}$ is a strong sm-uniqueness set for $\mathbb{R}^{d}$.
This Theorem allows us to easily generate strong sm-uniqueness sets, by uniformly sampling $k$ random directions on the unit sphere. With these results in hand we are ready to study the two statistical problems mentioned above.
3 A new estimator for Gaussian mixtures
3.1 Introduction
In this section, we shall develop a new estimator of a multivariate mixture of Gaussian distributions based on two ideas: random projections to obtain one-dimensional mixtures of normal distributions, and one-dimensional estimators of the mixture based on characteristic functions (see [76] and [82] for the second idea). We shall compare our random-projection method (RP) with the expectation-maximization (EM) algorithm.
There is a large literature on this subject. See for instance the well-known work in [54], the recent work [27] and the references therein. The article [54] considers general Gaussian multivariate models being $\epsilon$-statistically learnable, that is, $F=\sum_{i=1}^{n}\lambda_{i}F_{i}$ satisfying $\min_{i}\lambda_{i}\geq\epsilon$ and $\min_{i\neq j}\frac{1}{2}\int|f_{i}-f_{j}|\geq\epsilon$. In [27], very strong results are shown for the location case, where the covariances are all equal and known. In contrast, we just prove strong consistency, but this is done for the general case where the covariances are unknown and may be different, in an almost universal framework assuming only a differentiability condition on the characteristic functions, stated before [34, Theorem 4.1] below.
3.2 Description of the algorithm
We propose a two-step sequential estimation procedure based on random projections and the empirical characteristic function (ECF).
Let $P$ be a Gaussian mixture on $\mathbb{R}^{d}$, with density
$\sum_{j=1}^{m}{\lambda_{j}}\frac{1}{(2\pi\det(\Sigma_{j}))^{1/2}}\exp\Bigl(-\tfrac{1}{2}(x-\mu_{j})^{T}\Sigma_{j}^{-1}(x-\mu_{j})\Bigr),\quad(x\in\mathbb{R}^{d}),$
where $\lambda_{1},\dots,\lambda_{m}\in[0,1]$, $\sum_{j=1}^{m}\lambda_{j}=1$, $\mu_{1},\dots,\mu_{m}\in\mathbb{R}^{d}$, and $\Sigma_{1},\dots,\Sigma_{m}$ are $d\times d$ positive-definite matrices. We denote $\Lambda:=(\lambda_{1},\dots,\lambda_{m})$, $\bm{\mu}:=(\mu_{1},\dots,\mu_{m})$, $\bm{\Sigma}:=(\Sigma_{1},\dots,\Sigma_{m})$, and $\Theta:=(\Lambda,\bm{\mu},\bm{\Sigma})$.
Given a unit vector $u\in\mathbb{R}^{d}$ and $\theta\in\Theta$, $P_{\langle u\rangle}(\theta)$ denotes the projection of the Gaussian mixture with parameters $\theta$ in the direction $u$. Given an iid sample $X_{1},\dots,X_{N}$, let $P_{N,\langle u\rangle}$ be the projection of the empirical measure in the same direction.
In the first step, the mixture weights $\Lambda$ are estimated by fitting univariate projected mixtures along several random directions. This exploits the fact that linear projections of a Gaussian mixture are again Gaussian mixtures, with transformed means and variances. The estimation is performed by minimizing the discrepancy between the empirical and theoretical characteristic functions of the projected data. Averaging across multiple random directions yields a global estimator of the weights.
In the second step, with the weights fixed, the projected means and variances are re–estimated. From these, the original mean vectors and covariance matrices of the mixture are reconstructed by solving quadratic optimization problems.
The overall procedure is summarized in Algorithm 1.
Algorithm 1 Two-step estimation via random projections and ECF with alignment
1:Sample $u_{1},\dots,u_{k}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$
2:Step 1: For each $u_{r}$, estimate $(\widehat{\upsilon}_{u_{r},(1)},\widehat{\sigma}^{2}_{u_{r},(1)},\widehat{\Lambda}_{u_{r},(1)})$ by minimizing $(Z_{u}-F_{u})^{\top}W(Z_{u}-F_{u})$
3:Alignment: Choose a pivot direction $u_{r^{\ast}}$ (e.g., with largest separation between $\widehat{\upsilon}_{u_{r},(1),j}$), order its components by $\widehat{\upsilon}_{u_{r^{\ast}},(1),j}$, and treat the ordered triples $(\widehat{\Lambda}_{u_{r^{\ast}},(1),j},\widehat{\upsilon}_{u_{r^{\ast}},(1),j},\widehat{\sigma}^{2}_{u_{r^{\ast}},(1),j})$ as prototypes. For each $u_{r}$, solve an assignment problem (Hungarian/global arg min) that matches the triples $(\widehat{\Lambda}_{u_{r},(1),j},\widehat{\upsilon}_{u_{r},(1),j},\widehat{\sigma}^{2}_{u_{r},(1),j})$ to the pivot prototypes, obtaining aligned parameters $(\widetilde{\upsilon}_{u_{r},(1)},\widetilde{\sigma}^{2}_{u_{r},(1)},\widetilde{\Lambda}_{u_{r},(1)})$.
4:Set $\widehat{\Lambda}_{(1)}=\frac{1}{k}\sum_{r=1}^{k}\widetilde{\Lambda}_{u_{r},(1)}$
5:Reconstruction: From $\{\widehat{\upsilon}_{u_{r},(2)},\widehat{\sigma}^{2}_{u_{r},(2)}\}$ solve, for $j=1,\dots,n$,
$\widehat{\mu}_{j}$ $:=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\Bigl(\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},(2),j}\Bigr)^{2},$ (4)
$\widehat{\Sigma}_{j}$ $:=\operatorname*{argmin}_{\Sigma_{j}\in S^{+,d}}\sum_{r=1}^{k}\Bigl(\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\Bigr)^{2},$ (5)
with $S^{+,d}$ the cone of $d\times d$ positive semidefinite matrices.
[…]
In Algorithm 1, the parameters are initialized separately in each random one-dimensional projection and the main driver of the initialization is a $k$-means split. Specifically, for a randomly sampled direction $u$ (normalized to unit length), the data are projected as $Y_{i}=u^{\top}X_{i}$ and we run $k$-means with two clusters on $\{Y_{i}\}_{i=1}^{N}$. The resulting two clusters provide a data-driven starting point: the initial projected locations are set to the within-cluster sample means, the projected scales to the within-cluster sample variances (truncated below by a small constant for numerical stability), and the initial mixing weight is taken as the empirical cluster proportion. To avoid poor starts, we only keep projections where both clusters contain at least a minimum number of observations and the two $k$-means centroids are sufficiently separated relative to $\mathrm{sd}(Y)$. After ordering components by increasing projected mean, the $k$-means-based initialization is optionally refined by a single soft-EM update for the mixing weight, and then passed to the GEL estimation step in that projection.
The reconstruction of each covariance matrix $\Sigma_{j}$ from the projected variances is obtained by minimizing the quadratic expression
$f(\Sigma)=\sum_{r=1}^{k}\bigl(u_{r}^{\top}\Sigma u_{r}-s_{r,j}\bigr)^{2},\qquad\Sigma\in\mathbb{S}^{d}.$
This objective is always convex in $\Sigma$, but uniqueness of the solution requires strict convexity, which in turn depends on the geometry of the directions $\{u_{r}\}$. The following proposition makes this condition explicit.
Proposition 1 (Uniqueness of covariance reconstruction).
Let $u_{1},\dots,u_{k}\in\mathbb{S}^{d-1}$ and consider
$f(\Sigma)=\sum_{r=1}^{k}\bigl(u_{r}^{\top}\Sigma u_{r}-s_{r}\bigr)^{2},\qquad\Sigma\in\mathbb{S}^{d}.$
If $\{u_{r}u_{r}^{\top}:r=1,\dots,k\}$ spans $\mathbb{S}^{d}$, equivalently if $\mathrm{rank}(A)=d(d+1)/2$ where the $r$-th row of $A$ is $\mathrm{vecsym}(u_{r}u_{r}^{\top})$, then $f$ is strictly convex on $\mathbb{S}^{d}$ and admits a unique minimizer. In particular, the reconstruction equations (4)(5) have a unique solution for each covariance matrix $\Sigma_{j}$.
Sketch of proof.
Identifying $\mathbb{S}^{d}$ with $\mathbb{R}^{m}$, $m=d(d+1)/2$, via $\theta=\mathrm{vecsym}(\Sigma)$, we can write
$f(\Sigma)=\bigl\|A\,\theta-s\bigr\|_{2}^{2},$
with $A\in\mathbb{R}^{k\times m}$ having $r$-th row $\mathrm{vecsym}(u_{r}u_{r}^{\top})$. The spanning condition $\{u_{r}u_{r}^{\top}\}$ spans $\mathbb{S}^{d}$ is equivalent to $\mathrm{rank}(A)=m$, so $A^{\top}A$ is positive definite. Hence the Hessian of $f$ with respect to $\theta$ is $2A^{\top}A\succ 0$, and $f$ is strictly convex, with a unique minimizer $\theta^{\star}$, hence a unique $\Sigma^{\star}$. ∎
Further details are provided in Appendix A.
Remarks.
(i) If $\Lambda$ is known or estimated by another mechanism, then the estimation procedure starts at Step 2. For example, as in [24], this happens when one starts with a cluster procedure that has already been performed. This allows to have a $\Lambda$-estimator across the group sizes obtained.
(ii) Robust estimators can be obtained by replacing the $L^{2}$-minimization by the $L^{1}$-one, that is, by solving the equations
$\widehat{\mu}_{j}=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\bigl|\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},j}\bigr|,\ \ \widehat{\Sigma}_{j}=\operatorname*{argmin}_{\Sigma_{j}\in S^{+,d}}\sum_{r=1}^{k}\bigl|\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},j}\bigr|,$
for all $j=1,2,\dots,n$.
Alternatively, in the spirit of the proposed median of means, see for instance [25] and the references therein, one can proceed as as follows. We split the sample into $L$ disjoint subsamples. For each subsample and $j$ we calculate $\widehat{\mu}_{j,\ell}$ and $\widehat{\Sigma}_{j,\ell}$, $j=1,\dots n$, $\ell=1,\ldots,L$ according to equations (4) and (5), and we take the median of the corresponding $L$ values obtained for each $j$. Then we proceed as before.
(iii) The $\operatorname*{argmin}$ in (4) can be easily solved using least squares. To solve the constrained optimization problem of estimating covariance matrices under positive semi-definiteness, see (5), we apply a primal-dual interior-point method (IPM). This approach is well-suited for semidefinite programming (SDP), as it ensures convergence to the global optimum within the cone of symmetric positive semidefinite matrices. The method combines Newtonian steps with a logarithmic barrier function to keep iterates in the feasible region, which in our case corresponds to the set $\mathbb{S}^{+,d}$. This strategy is particularly appropriate given the convexity of our loss function and the structure of the constraint. For background and theoretical guarantees, see [81, 57, 68].
Since $f(\Sigma)=\Bigl(\langle u_{r},\Sigma u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\Bigr)^{2}$ is convex and continuous on $\mathbb{S}^{+,d}$, the optimization problem admits a unique solution. Consequently, the penalized objective function used in the IPM Algorithm, $f_{t}(\Sigma)=tf(\Sigma)-\log\det(\Sigma)$, attains a unique minimum over $\mathbb{S}^{+,d}$; see, e.g., [7, 10].
(iv) While semidefinite programming (SDP) is computationally efficient for small-to-moderate dimensions, various alternatives are available for high-dimensional settings. Examples include first-order methods such as the Alternating Direction Method of Multipliers (ADMM), low-rank matrix approximations when the covariance structure allows, and the exploitation of sparsity or block structure in the projection matrices to reduce computational cost, see for example [47] and [79].
(v) In general, the problem becomes increasingly challenging as the dimension of the space grows. The parameter space expands substantially when allowing for general covariance matrices, which increases complexity and renders the algorithm more unstable, since one must optimize over symmetric positive definite matrices in addition to the weights and means. Recently, in [83], a novel and very general problem is studied that includes the more complex setup of compound probability distributions under the restriction of identical and isotropic covariances, which for our setting requires that all covariance matrices be the identity.
(vi) We note that the algorithm can also be applied to mixtures of multivariate $t$-distributions.
(vii) It is important to note that Step 1 can be implemented with any other consistent univariate-mixture estimators at each direction $u_{r}$, for instance the univariate EM algorithm, instead of the ECF-based criterion.
3.3 A consistency result
The results in [76] and [82] that we will use below are based on a general theorem in [43], where the only assumption is that the characteristic function is differentiable. More precisely, it is assumed that the regular case holds, namely that $I_{n}(t)$ can be differentiated under the integral sign. Proposition 1 in [76], which is based on the work in [31], [32], can be stated as follows.
Proposition 2.
Let $\{t_{1},\ldots,t_{m}\}$ be distinct fixed grid points. Then the estimator $\widehat{\theta}$ of $\theta$ is strongly consistent and asymptotically normally distributed, with a given covariance matrix.
How to choose the sequence $\{t_{1},\ldots,t_{m}\}$ as well as the size $m$ in an optimal way for a given sample size $N$ is still an open problem. However, in [32] it is suggested that the sequence should be taken equally spaced, namely $t_{j}=\tau j,\ j=1,\ldots m$ for $\tau\in\mathbb{R}^{+}$.
Theorem 7.
Under the weak assumption that,
$I^{\prime}_{n}(\theta)=\int\frac{\partial}{\partial\theta}|\widehat{\psi}_{n}(t)-\psi(t,\theta)|^{2}dP(t),$
the estimators $\widehat{\Lambda}_{(1)},\widehat{\mu}_{j},\widehat{\Sigma}_{j}$ for $j=1,\dots,n$ are strongly consistent.
Proof.
First observe that, from the results in [76] and [82] for the finite set of estimators for the projected one-dimensional projections, we can derive the strong consistency of $\widehat{\Lambda}_{(1)}$. See [76, Proposition 1], where strong consistency and asymptotic normality are established.
For the next step, using the value of $\widehat{\Lambda}_{(1)}$, we consider the estimators of $\widehat{\mu}_{j}$ and $\widehat{\Sigma}_{j}$ for $j=1,\dots n$ given by the least-squares equations (4) and (5). Since
$\sum_{r=1}^{k}\bigl(\langle u_{r},\mu_{j}\rangle-\upsilon_{u_{r},j}\bigr)^{2}=0\quad\text{and}\quad\sum_{r=1}^{k}\bigl(\langle u_{r},\Sigma_{j}u_{r}\rangle-\sigma^{2}_{u_{r},j}\bigr)^{2}=0$
for the true values of $\mu_{j}$ and $\Sigma_{j}~(j=1,\dots n)$, we have
$\widehat{\mu}_{j}=\operatorname*{argmin}_{\mu_{j}}\sum_{r=1}^{k}\Bigl(\bigl(\langle u_{r},\mu_{j}\rangle-\widehat{\upsilon}_{u_{r},j}\bigr)^{2}-\bigl(\langle u_{r},\mu_{j}\rangle-\upsilon_{u_{r},j}\bigr)^{2}\Bigr)\to 0\quad\text{a.s.}$
and
$\widehat{\Sigma}_{j}=\operatorname*{argmin}_{\Sigma_{j}\in S^{+,d}}\sum_{r=1}^{k}\Bigl(\bigl(\langle u_{r},\Sigma_{j}u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},j}\bigr)^{2}-\bigl(\langle u_{r},\Sigma_{j}u_{r}\rangle-\sigma^{2}_{u_{r},j}\bigr)^{2}\Bigr)\to 0\quad\text{a.s.}$
for all $j=1,\dots,n$. This concludes the proof. ∎
The only assumption in our Theorem 7 is that $I_{n}(t)$ can be differentiated under the integral sign. As mentioned in [43, p. 258], the only random variables excluded are those with characteristic functions vanishing outside an interval depending on the parameter $\theta$ and those for which $\psi$ is not differentiable with respect to $\theta$. These conditions are fulfilled for the univariate mixtures that we have under consideration. Indeed, the differentiation condition follows for instance from [63].
Therefore our strong consistency result is universal, i.e., it holds without further assumptions for Gaussian or $t$-mixtures. Moreover, the theorem-level requirement in [34] shows that it suffices to take
$k\;\geq\;\frac{1}{2}(2m-1)(d^{2}+d-2)+1$
projection directions, where $m$ denotes the number of mixture components and $d$ the ambient dimension. Hence the required number of directions grows linearly in $m$ and quadratically in $d$.
No restriction is placed on the sample size.
An interesting related but different approach is given by Moitra and Valiant in [55] for the estimation of the parameters of mixture of multivariate Gaussians (GMM), where they provide an algorithm that has a polynomial running time. To state their main result, we provide some definitions. In what follows, we write $k]:=\{1,2,\dots,k\}$.
Definition 2.
(i) Given two $d$-dimensional GMMs of $m$ Gaussians, $F=\sum_{i}w_{i}N(\mu_{i},\Sigma_{i})$ and $\widehat{F}=\sum_{i}\widehat{w}_{i}N(\widehat{\mu}_{i},\widehat{\Sigma}_{i})$, we call $\widehat{F}$ an $\epsilon$-close estimate for $F$ if there is a permutation function $\pi:[k]\to[k]$ such that, for all $i\in[k]$,
$\|w_{i}-\widehat{w}_{\pi[i]}\|\leq\epsilon\quad\text{and}\quad D(N(\mu_{i},\Sigma_{i}),N(\widehat{\mu}_{\pi(i)},\widehat{\Sigma}_{\pi(i)}))\leq\epsilon.$
(ii) We call a GMM $F=\sum_{i}w_{i}F_{i}$ $\epsilon$-statistically learnable if $\min_{i}w_{i}\geq\epsilon$ and $\min_{i}D(F_{i},F_{j})\geq\epsilon$.
The result of Moitra and Valiant is given in the following theorem.
Theorem 8 ([55, Theorem 1]).
Given an $d$-dimensional mixture of $m$ Gaussians $F$ that is $\epsilon$-statistically learnable, there is an algorithm that, with probability at least $1-\delta$, outputs an $\epsilon$-close estimate $\widehat{F}$, and the running-time and data requirements of the algorithm (for any fixed $m$) are polynomial in $n$, $1/\epsilon$ and $1/\delta$.
If one is only interested on D(F,F^)D($\widehat{F}$), then in [55, Corollary 2] there is stated a similar result without the restriction that $F$ be $\epsilon$-statistically learnable.
Sharp results for the estimation of univariate finite mixtures are also provided by Heinrich and Kahn in [44], where they study the rates of convergence of the parameters of the mixture, and, under some regularity and strong identifiability conditions, find the optimal local minimax rate of estimation of a univariate mixture with $m$ components. However they do not address the multivariate case.
3.4 Simulations
3.4.1 Example 1.
Using the above algorithm, we estimate the parameters of a mixture of two bivariate $t$-distributions $t(\nu,\mu,\Sigma)$, where $\nu$, $\mu$, and $\Sigma$ denote the degrees of freedom, location, and scale matrix parameters, respectively. Consider
$F$ $=\lambda_{1}\,t(\nu,\mu_{1},\Sigma_{1})+(1-\lambda_{1})\,t(\nu,\mu_{2},\Sigma_{2}),$
$\mu_{1}$ $=(0,0),\quad\mu_{2}=(\eta,0),$
$\Sigma_{1}$ $=\begin{pmatrix}1&0\\ 0&1/2\end{pmatrix},\quad\Sigma_{2}=\begin{pmatrix}1/2&0\\ 0&1\end{pmatrix}$,
[…]
Based on $200$ i.i.d. draws from $F$ (see Figure 1), our goal is to estimate $\mu_{1}$, $\mu_{2}$, $\Sigma_{1}$, and $\Sigma_{2}$. We implement the proposed algorithm by projecting the data onto $k=50$ randomly selected directions. This experiment is replicated $100$ times. In addition, we consider four separability scenarios by letting $\eta\in\{1/2,\,1,\,3/2,\,2\}$, which progressively increases the separation between the component locations.
Fig. 1: Bivariate $t$-mixture samples for increasing separation $\eta$ (four panels: $\eta\in\{1/2,1,3/2,2\}$).
In this Example, we present a small comparative study of the estimation derived by the random projections method (RP) with respect to the expectation-maximization method for mixtures of t-distributions (we denote EM-st). This last method, developed in [62], is implemented by means of the function fmmst of the EMMIXuskew package of the R language. In this case the skew parameter was prefixed to 0.
Table 1 reports the average parameter estimates across the replicates, together with their corresponding standard deviations. In addition, we report the (replicate-averaged) confusion matrices (see Table 2) computed over the $100$ Monte-Carlo runs—comparing the true component labels with the posterior (MAP) allocations obtained after parameter estimation, for both the EM-st and the RP procedures. These confusion matrices are provided in Appendix B.
In each replicate, we compute estimation errors by comparing the estimated parameters with their true values. Specifically, we construct boxplots of (i) the $L^{2}$-error for the mixing weight $\lambda_{1}$, (ii) the $L^{2}$-error for the mean vectors, and (iii) the Frobenius-distance for the variance–covariance matrices. These boxplots are reported for each separability scenario $\eta\in\{1/2,1,3/2,2\}$; see Figure 3 for EM-st and RP, respectively. Overall, both algorithms exhibit comparable performance across scenarios. In Figure 2, the boxplots show that the estimation errors for the mixing coefficients $\lambda$ are similar for both methods.
Fig. 2: Mixing-weight error for $\lambda_{1}$: EM-st vs RP across separability scenarios $\eta\in\{1/2,1,3/2,2\}$ (boxplots over 100 replicates).
Fig. 3: Comparison between the EM-st (EM) and RP methods via boxplots of estimation errors across the four separability scenarios $\eta\in\{1/2,1,3/2,2\}$. Errors are measured using the $L^{2}$-distance for the mean vectors, and the Frobenius distance for the variance–covariance matrices.
3.4.2 Example 2
We repeat the previous simulation setup, but now the observations are generated from a two-component bivariate Student $t$-mixture with $\nu=4$ degrees of freedom, contaminated by a fraction $\gamma$ of uniform noise on an axis-aligned square. Specifically, for each replicate we generate $N=500$ observations from
$F=(1-\gamma)\Big\{\lambda_{1}\,t_{\nu}(\mu_{1},\Sigma_{1})+\lambda_{2}\,t_{\nu}(\mu_{2},\Sigma_{2})\Big\}+\gamma\,\mathcal{U}\!\big([0,4]\times[0,4]\big),$
where $\lambda_{1}=0.3$ and $\lambda_{2}=0.7$, and
$\mu_{1}=(0,0),\qquad\mu_{2}=(2,0),$
$S_{1}=\begin{pmatrix}1&0\\[1.00006pt] 0&1/2\end{pmatrix},\qquad S_{2}=\begin{pmatrix}1/2&0\\[1.00006pt] 0&1\end{pmatrix}$.
The contamination distribution is uniform on the square centered at $(2,2)$ with side length $4$, i.e., $\mathcal{U}([0,4]\times[0,4])$. The parameter $\gamma$ controls the proportion of outliers, and we consider $\gamma\in\{0.05,\,0.10,\,0.15\}$. Figure 4 displays one simulated dataset for each noise scenario considered in our study. Each panel corresponds to a different contamination level $\gamma\in\{0.05,0.10,0.15\}$ and illustrates a single realization from the bivariate Student $t$-mixture with uniform outliers; hence, the figure provides a visual summary of how the amount of uniform noise increases across scenarios.
Fig. 4: Simulated bivariate Student $t$-mixture with uniform contamination. Each panel shows a sample of size $N=500$ from a two-component $t_{\eta}$ mixture with $\eta=4$, mixing weights $(\lambda_{1},\lambda_{2})=(0.3,0.7)$, locations $\mu_{1}=(0,0)$ and $\mu_{2}=(2,0)$, and scatter matrices $\Sigma_{1}=\mathrm{diag}(1,1/2)$ and $\Sigma_{2}=\mathrm{diag}(1/2,1)$. A proportion $\gamma\in\{0.05,0.10,0.15\}$ of observations is replaced by outliers drawn uniformly from $0,4]\times[0,4]$.
In this case we re-estimate the original mixture parameters, and the $L^{2}$-errors are plotted in Figure 5.
In this framework, our method is compared to a mixture estimation using a robust variant of EM (which we denote by RobEM). These estimators are introduced in [39] and are implemented in the RGMM package of the R language. The methodology they propose consists of modifying the EM algorithm, in particular the M-step, where they replace the mean and variance estimates by robust versions derived from the median and the median of the covariance matrix, respectively.
In the robust contamination setting, the RP approach exhibits an overall better performance than the robust EM alternative. This improvement is particularly clear for the mixing proportion parameter $\lambda$, whose estimation is more stable under RP across the noise scenarios, as well as for the scatter parameters: the RP reconstructions of the variance–covariance (scatter) matrices show smaller discrepancies with respect to the true matrices, indicating a higher robustness of RP to the uniform outliers when recovering second-order structure, see Figure 5 and Figure 6.
Fig. 5: Comparison of estimation errors for the mixing proportion $\lambda_{1}$ under uniform contamination. Each panel corresponds to a different outlier proportion $\gamma\in\{0.05,0.10,0.15\}$ (left to right), and boxplots summarize the errors over $100$ Monte Carlo replicates for the RobEM and RP methods.
Fig. 6: Comparison of estimation errors for the remaining parameters under uniform contamination. Each panel corresponds to a different outlier proportion $\gamma\in\{0.05,0.10,0.15\}$ (left to right), and boxplots summarize the errors over $100$ Monte Carlo replicates for the RobEM and RP methods. Errors are measured using the $L^{2}$-distance for the mean vectors and the Frobenius distance for the variance–covariance matrices.
Table 3 reports the average parameter estimates across the replicates, together with their corresponding standard deviations, under the contamination setting with uniform outliers. In addition, we report the (replicate-averaged) confusion matrices (see Table 4) computed over the $100$ Monte-Carlo runs—comparing the true component labels with the posterior (MAP) allocations obtained after parameter estimation, for both the robust EM and the RP procedures, see Appendix B.
3.5 An example with real data
The relationship between school performance and students’ socioeconomic and cultural status has been widely studied, see [80]. This association is even stronger in the Latin American countries, see [28].
In Uruguay, this trend can be seen in the educational evaluation studies of the National Institute for Educational Evaluation (INEEd). A representative sample of $6437$ students in the third year of secondary education in the country in 2022 is considered. The sample is comprised of students attending two types of educational institutions: Public (free and funded by the state) and Private (depending on the payment of tuition by their students or private sources of financing). In the sample, $4852$ and $1585$ students attend public and private schools, respectively. By means of a multiple-choice test, using Item Response Theory, a score in Mathematics is assigned to the item. In addition, an index of each student’s socioeconomic and cultural level is constructed from data collected in a personal questionnaire. [^2] The database and indexes mentioned above are open and available at https://www.ineed.edu.uy/nuestro-trabajo/aristas/
Figure 7 depicts a bivariate histogram in hexagonal cells. This Figure shows the positive relationship between the two measures, where an asymmetric distribution is observed.
Fig. 7: Hexagonal binning for bivariate data. The plane is tessellated by a regular grid of hexagons (cells). Then the hexagons are plotted using an intensity of color proportional to the number of points falling in each cell. Higher socio-economic level for smaller values on the $x$-axis, better mathematical ability for larger values on the $y$-axis.
The objective is to model this bivariate distribution by means of a mixture of normals and to explain a possible reason for this asymmetry. Using the method that we have developed, the following estimated parameters of a mixture of two Gaussians are obtained: $\widehat{\lambda}_{1}=0.75,\ \widehat{\mu}_{1}=(-0.14,-0.15),\ \widehat{\mu}_{2}=(1.16,0.47)$, and
$\widehat{\Sigma}_{1}=\begin{pmatrix}0.39&0.15\\ 0.15&0.65\end{pmatrix},\quad\widehat{\Sigma}_{2}=\begin{pmatrix}0.16&0.08\\ 0.08&0.48\end{pmatrix}$.
These two data patterns are possibly explained by the type of institution that the students attend. In this example, we consider $\lambda$ to be known a priori, i.e., the proportion of students attending Public educational institution is $0.75$. If we calculate the vector of means and the matrix of empirical variances and covariances of the sample conditioned to each institution, we obtain $\widehat{\mu}_{\text{Public}}=(-0.16,-0.180),\ \widehat{\mu}_{\text{Private}}=(1.18,0.51)$, and
$\widehat{\Sigma}_{\text{Public}}=\begin{pmatrix}0.45&0.10\\ 0.10&0.75\end{pmatrix},\quad\widehat{\Sigma}_{\text{Private}}=\begin{pmatrix}0.20&0.12\\ 0.12&0.59\end{pmatrix}$.
These values are similar to those obtained in the mixture estimation.
An important aspect of this example is the comparison of student performance in acquired mathematical skills between those attending private schools and those attending public schools. The comparison considers two variables: the score assigned to the item and the socioeconomic and cultural level of the child’s home. Comparing the means $\widehat{\mu}_{1}=(-0.14,-0.15)$ for those attending public schools with $\widehat{\mu}_{2}=(1.16,0.47)$ for those attending private schools, we see a clear drop in the performance of students attending public schools, particularly marked in the first coordinate, but also significant in the second. This highlights a problem to be considered by the country’s policymakers.
4 Comparing random partitions by model-based methods
4.1 Introduction
The problem of comparing two different partitions of a finite set of observations is an important classical problem in the clustering literature. There is a large amount of work on the subject with several different proposals. Our aim in this section is to propose a new approach to this problem. We consider not only the cluster labels assigned by the clustering algorithm but also the distribution of the resulting groups. Note that we do not address the general problem of model-based clustering, which is a broad topic (see [37] and [9]).
The most popular measures are given by the Rand index [66], the adjusted Rand index (Hubert and Arabie [45]), the Jaccard index, the Dunn index, the silhouette index, and the Xie–Beni index among others. See for instance the reviews by Arabie and Boorman [3], Fowlkes and Mallows [33] among others. More recently Youness and Saporta [84] proposed a different approach to the problem. Most of these existing indices lack real mathematical analysis, and almost no information exists about their distribution, see [53]. Also, the problem becomes harder for high-dimensional data. Our model-based approach provides a good interpretation of this measure in terms of the underlying unknown distributions.
On the other hand, a probabilistic model used quite often in Bayesian statistics as well in cluster analysis is to consider that the data are generated by a mixture of multivariate normal distributions. Indeed, a popular and well-studied clustering method is to use Gaussian mixture models (GMM) for data clustering (model-based clustering). One can perform hard clustering or soft clustering. For hard clustering the GMM model assigns each datum to the component that maximizes the component posterior probability, given the datum, and each datum is assigned to only one cluster. Soft clustering assigns each instance a probability of belonging to a cluster. See for instance Gormley et al. [40] for details.
From the recent article [40] we quote: “Through its basis in a statistical modeling framework, model-based clustering provides a principled and reproducible approach to clustering. In contrast to heuristic approaches, model-based clustering allows for robust approaches to parameter estimation and objective inference on the number of clusters, while providing a clustering solution that accounts for uncertainty in cluster membership …”
Instead of measuring agreement between two different partitions, we consider the problem of measuring the agreement between the probability measures that have produced those partitions (the output of two different model-based clustering procedures), when both distributions correspond to finite mixtures of multivariate normal or $t$-distributions. In particular, we are able to treat high-dimensional data.
Our proposal does not, in principle, provide distances between partitions directly. Instead, it provides a measure between two model-based distributions associated with the partitions. More precisely, in our framework, the objects being compared are the mixture distributions associated with the two partitions (i.e., the distributions induced by the model-based clusterings), instead of the label assignments themselves. For this reason, we report $D_{k}$ and $\mathrm{MA}_{k}$ defined below as descriptive measures of agreement between partitions.
As pointed in Baudry et al. [5], several problems appear when performing Gaussian-model-based clustering, affecting the number of clusters obtained by BIC. Some nice solutions are proposed. However these problems do not affect us. An important aspect of our approach is that we do not need to perform the clustering method to calculate the measuring agreement, and that also we can easily deal with high-dimensional data. In other words, the method does not require a prior criterion for selecting the number of clusters, and therefore avoids the issues associated with poor estimation of the number of clusters.
4.2 Description of the algorithm
Suppose that we want to compare the results obtained from two clustering procedures, both applied to the same set of variables. We consider two independent samples $\aleph_{1}=\{X_{1},\dots,X_{\ell}\}\subset\mathbb{R}^{d}$ and $\aleph_{2}=\{Y_{1},\dots,Y_{r}\}\subset\mathbb{R}^{d}$, with distributions $P$ and $Q$ respectively, both being mixtures of $m$ multivariate Gaussian or $t$-distributions.
Our method relies on random projections and Kolmogorov–Smirnov (KS) distances. Specifically, each sample is projected onto several random one-dimensional subspaces, empirical distributions are computed, and the KS distances are evaluated between the projected samples. Aggregating these distances provides a measure of discrepancy between the two clustering results.
We consider two iid samples
$\aleph_{1}=\{X_{1},\dots,X_{\ell}\}\subset\mathbb{R}^{d}\ \mbox{and}\ \aleph_{2}=\{Y_{1},\dots,Y_{r}\}\subset\mathbb{R}^{d},$
drawn from distributions $P$ and $Q$, each being a mixture of $m$ multivariate Gaussian or $t$-distributions. Let $P_{\ell}$ and $Q_{r}$ be the empirical distributions of $\aleph_{1}$ and $\aleph_{2}$. Applying Theorems 4 and 5, the discrepancy between the clusters is measured by the average of the projected KS-distances between $P_{\ell}$ and $Q_{r}$ along a strong sm-uniqueness set of random directions (Theorem 6).
For each direction $u_{i}$, denote the projected samples by
$\mathcal{M}_{1i}=\{\langle u_{i},X_{j}\rangle\}_{j=1}^{\ell},\qquad\mathcal{M}_{2i}=\{\langle u_{i},Y_{j}\rangle\}_{j=1}^{r},$
with corresponding empirical CDFs $F_{1i}$ and $F_{2i}$. The one-dimensional Kolmogorov–Smirnov statistic is
$\mathrm{KS}(i)=\sup_{t\in\mathbb{R}}|F_{1i}(t)-F_{2i}(t)|.$
We then define
$D_{k}:=\max_{1\leq i\leq k}\mathrm{KS}(i).$
As $\min(\ell,r)\to\infty$, the asymptotic distribution of $D_{k}$ satisfies
$\lim_{\min(\ell,r)\to\infty}\Pr(D_{k}\leq t)=\lim_{\min(\ell,r)\to\infty}\Big[\Pr(\mathrm{KS}(1)\leq t)\Big]^{k},$
which provides the basis for an asymptotic test. A drawback is the loss of the distribution-free property, as the distribution of $D_{k}$ depends on the covariance structure of the underlying processes.
The overall scheme is summarized in Algorithm 2.
Algorithm 2 Agreement measure via random projections and Kolmogorov–Smirnov distances
1:Integers $m,d$; samples $\aleph_{1}=\{X_{1},\dots,X_{\ell}\}\subset\mathbb{R}^{d}$, $\aleph_{2}=\{Y_{1},\dots,Y_{r}\}\subset\mathbb{R}^{d}$; fix $k\geq\tfrac{1}{2}(2m-1)(d^{2}+d-2)+1$
2:Draw $k$ random directions $u_{1},\dots,u_{k}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$
3:for $i=1$ to $k$ do
4:  Project samples: $\mathcal{M}_{1i}=\{\langle u_{i},X_{j}\rangle\}_{j=1}^{\ell}$,  $\mathcal{M}_{2i}=\{\langle u_{i},Y_{j}\rangle\}_{j=1}^{r}$
5:  Compute empirical CDFs $F_{1i},F_{2i}$ from $\mathcal{M}_{1i}$ and $\mathcal{M}_{2i}$
6:  $\mathrm{KS}(i)\leftarrow\sup_{t\in\mathbb{R}}|F_{1i}(t)-F_{2i}(t)|$
7:end for
8:return $\mathrm{D}_{k}\leftarrow\max_{i=1,\dots,k}\mathrm{KS}(i)$
Remarks.
(i) An alternative measure is the average
$\mathrm{MA}_{k}:=\frac{1}{k}\sum_{i=1}^{k}\mathrm{KS}(i),$
where $\mathrm{KS}(i)$ is the KS distance between the empirical distributions of $\mathcal{M}_{1i}$ and $\mathcal{M}_{2i}$. If $F_{i}$ and $G_{i}$ are the true projected distributions, then by the triangle inequality,
$0\leq\mathrm{KS}(i)\leq\mathrm{KS}(F_{\mathcal{M}_{1i}},F_{i})+\mathrm{KS}(F_{i},G_{i})+\mathrm{KS}(G_{\mathcal{M}_{2i}},G_{i}),$
which implies
$\lim_{\min(\ell,r)\to\infty}\mathrm{MA}_{k}=\frac{1}{k}\sum_{i=1}^{k}\mathrm{KS}(F_{i},G_{i}).$
Theorems 4 and 5 show that if this limit is zero, then $P=Q$. However, the asymptotic distribution of $\mathrm{MA}_{k}$ is unknown, since projections onto different directions are only conditionally independent.
(ii) If the sample sizes $\ell$ and $r$ are large, one may split them into $k$ disjoint subsamples and use different subsamples for each direction $u_{i}$. In this case, $\mathrm{MA}_{k}$ (or $D_{k}$) becomes an average (or maximum) of $k$ independent one-dimensional KS statistics, which is distribution-free. This strategy, however, comes at the cost of reduced statistical power.
(iii) Pre-whitening via $\Sigma^{-1/2}$ is a convenient option to mitigate scale (and unit-of-measure) effects when the practitioner wishes to control for them; in that case, one may apply a whitening transform to the data before computing the projected KS aggregations.
4.3 Simulations
Consider two mixtures of two bivariate Gaussians, with parameters
$F_{1}$ $:=\lambda_{1}N\left(\mu_{1,1},\Sigma_{1,1}\right)+(1-\lambda_{1})N\left(\mu_{2,1},\Sigma_{2,1}\right),$
$F_{2}$ $:=\lambda_{2}N\left(\mu_{1,2},\Sigma_{1,2}\right)+(1-\lambda_{2})N\left(\mu_{2,2},\Sigma_{2,2}\right),$
where
$\mu_{1,1}$ $=\mu_{1,2}:=(1,-1),\mu_{2,1}:=(-2,2),\lambda_{1}=0.5,$
$\Sigma_{1,1}$ $=\Sigma_{1,2}:=\begin{pmatrix}1&0\\ 0&2\end{pmatrix}\quad\text{and}\quad\Sigma_{2,1}:=\begin{pmatrix}3&1\\ 1&4\end{pmatrix}$,
and where, given $\eta_{1},\eta_{2},\eta_{3}$, we set $\mu_{2,2}:=\mu_{2,1}+(\eta_{1},\eta_{1})$, $\Sigma_{2,2}:=(1+\eta_{2})\Sigma_{2,1}$, and $\lambda_{2}:=0.5+\eta_{3}$.
The distances between the two mixtures $F_{1}$ and $F_{2}$ are calculated for three scenarios:
Scenario 1:
$\eta_{1}$ on a grid of values in $[0,1]$ and $\eta_{2}=\eta_{3}=0$,
Scenario 2:
$\eta_{2}$ on a grid of values in $[0,2]$ and $\eta_{1}=\eta_{3}=0$,
Scenario 3:
$\eta_{3}$ on a grid of values in $[0,0.5]$ and $\eta_{1}=\eta_{2}=0$.
Fig. 8: Representative realization in each scenario comparing mixtures $F_{1}$ and $F_{2}$: Scenario 1 varies the location shift $\eta_{1}$ (with $\eta_{2}=\eta_{3}=0$), Scenario 2 varies the scale factor $\eta_{2}$ (with $\eta_{1}=\eta_{3}=0$), and Scenario 3 varies the mixing-weight perturbation $\eta_{3}$ (with $\eta_{1}=\eta_{2}=0$).
(a) Scenario 1
(b) Scenario 2
(c) Scenario 3
Fig. 9: Agreement between mixtures $F_{1}$ and $F_{2}$ using random projections: boxplots of $D_{k}=\max_{1\leq i\leq k}\mathrm{KS}(i)$ (left) and $\mathrm{MA}_{k}=\frac{1}{k}\sum_{i=1}^{k}\mathrm{KS}(i)$ (right), over 500 replicates with $k=100$ directions and 500 draws from each mixture per replicate. Scenario 1 varies location ($\eta_{1}$), Scenario 2 varies scale ($\eta_{2}$), and Scenario 3 varies mixing weights ($\eta_{3}$).
In Figure 8, for improved visualization, one representative simulation per scenario is displayed, corresponding to $\eta_{1}=1$, $\eta_{2}=2$, and $\eta_{3}=1/2$ in Scenarios 1, 2, and 3, respectively. The maximum aggregation measure $D_{k}$ quantifies the largest observed separation between the mixtures $F_{1}$ and $F_{2}$ across the $k$ projected KS comparisons. This measure is particularly sensitive to pronounced discrepancies (whether in location, dispersion, or component weights), thereby providing a robust indicator of distinct distributional features.
Figure 9 presents, for each scenario, boxplots of the $500$ replicates of $D_{k}$, obtained by simulating 500 observations from each mixture and repeating the procedure 500 times. When $\eta_{1}=\eta_{2}=\eta_{3}=0$, the mixtures coincide and $D_{k}$ attains minimal values. As the mixtures diverge, $D_{k}$ exhibits a clear increasing trend, faithfully reflecting the growing separation.
The same analysis was carried out using the average aggregation measure $\textit{MA}_{k}$, obtaining similar results, as shown in the right column of Figure 9.
5 Conclusions and future work
The article begins with a review of random-projection methods in statistics.
Then we consider two important statistical problems: that of estimating for mixtures of multivariate normal distributions and mixtures of $t$-distributions, and that of measuring the agreement between two different random partitions for cluster analysis.
We propose a new estimator for multivariate mixture distributions based on one-dimensional projections and their respective characteristic functions. We derive the strong consistency of them and, in a small simulation study, we compare with the more classical EM-type estimators, where we observe a very good behaviour.
A model-based measure of agreement is introduced and compared with more classical measures. Instead of looking at the number of matchings, we look at the distance between the mixture distributions associated with them, which provide much more information about the agreement between the partitions.
The relationship between school performance and students’ socioeconomic and cultural status is analyzed in a real-data example. We find two different patterns that are possible, explained by the type of institution that the students attend : public or private.
The number of directions $k$ into which we project is not a smoothing parameter that should be chosen in an optimal way. In practice we suggest to use a larger value for $k$ (without increasing too much the computational time) than the one provided by Theorem 4, which corresponds to the case when the mixtures are known, whereas we are dealing with empirical distributions which are close to the unknown underlying distributions but not exactly equal. Depending on the problem, among the directions chosen at random it may appear some directions which are close to bad ones. Increasing the value of $k$ solves the problem. In our experience after a reasonable value the performance stabilizes.
The two parts of the manuscript are linked by a common projection-based representation for Gaussian/$t$-mixtures: the same finite family of directions that supports identifiability and reconstruction in Section 3 also yields a natural distributional notion of partition agreement in Section 4, obtained by aggregating one-dimensional discrepancies across directions.
Some more computationally oriented work, which includes different mixtures of distributions, sample sizes and number of clusters, has been omitted from the present manuscript. Also, rates of convergence and the asymptotic distribution of the estimators remain as interesting open problems to be addressed in future work.
Acknowledgments
We are grateful to the four anonymous referees for their thoughtful suggestions and detailed feedback, which led to substantial improvements in the presentation and overall quality of the manuscript.
Fraiman and Moreno were supported by grant FCE-3-2022-1-172289 from ANII (Uruguay), the Program for the Development of Basic Sciences–PEDECIBA (Uruguay), and 22520220100031UD from CSIC (Uruguay). Ransford was supported by NSERC Discovery Grant RGPIN-2020-04263.
Appendix A Algorithm 1: Methodological aspects
This appendix details the implementation of Algorithm 1.
A.1 Step 1: Estimation of the mixing weights via one-dimensional ECF fits
Let $P$ be an $m$-component Gaussian mixture on $\mathbb{R}^{d}$ with parameters $\Theta=(\Lambda,\bm{\mu},\bm{\Sigma})$, where $\Lambda=(\lambda_{1},\dots,\lambda_{m})$, $\sum_{j=1}^{m}\lambda_{j}=1$, $\bm{\mu}=(\mu_{1},\dots,\mu_{m})$, and $\bm{\Sigma}=(\Sigma_{1},\dots,\Sigma_{m})$.
Fix a unit direction $u\in\mathbb{S}^{d-1}$ and define the projected data
$Y_{i}(u):=\langle u,X_{i}\rangle,\qquad i=1,\dots,N.$
Under the model, $Y_{i}(u)$ follows a univariate Gaussian mixture with component-specific projected means and variances
$\upsilon_{u,j}:=\langle u,\mu_{j}\rangle,\qquad\sigma^{2}_{u,j}:=\langle u,\Sigma_{j}u\rangle,\qquad j=1,\dots,m,$
that is,
$P_{\langle u\rangle}=\sum_{j=1}^{m}\lambda_{j}\,\mathcal{N}\!\big(\upsilon_{u,j},\sigma^{2}_{u,j}\big).$
Characteristic functions.
The (model) characteristic function of the projected mixture is
$\psi_{u}(t;\upsilon_{u},\sigma_{u}^{2},\Lambda):=\sum_{j=1}^{m}\lambda_{j}\exp\!\left(it\,\upsilon_{u,j}-\tfrac{1}{2}t^{2}\,\sigma^{2}_{u,j}\right),$
where $\upsilon_{u}=(\upsilon_{u,1},\dots,\upsilon_{u,m})$ and $\sigma_{u}^{2}=(\sigma^{2}_{u,1},\dots,\sigma^{2}_{u,m})$. Its empirical counterpart (ECF) is
$\widehat{\psi}_{u,N}(t):=\frac{1}{N}\sum_{i=1}^{N}\exp\!\big(it\,Y_{i}(u)\big)=\frac{1}{N}\sum_{i=1}^{N}\exp\!\big(it\,\langle u,X_{i}\rangle\big).$
Grid and moment vector.
Fix grid points $t_{1},\dots,t_{M}$ (e.g., $t_{\ell}=\tau\,\ell$ as in [32]). Define the empirical moment vector
$\widehat{Z}_{u}:=\big(\operatorname{Re}\widehat{\psi}_{u,N}(t_{1}),\dots,\operatorname{Re}\widehat{\psi}_{u,N}(t_{M}),\ \operatorname{Im}\widehat{\psi}_{u,N}(t_{1}),\dots,\operatorname{Im}\widehat{\psi}_{u,N}(t_{M})\big)^{\top},$
and the model moment vector
$Z_{u}$ $(\upsilon_{u},\sigma_{u}^{2},\Lambda)$
$:=\big(\operatorname{Re}\psi_{u}(t_{1};\upsilon_{u},\sigma_{u}^{2},\Lambda),\dots,\operatorname{Re}\psi_{u}(t_{M};\cdot),\ \operatorname{Im}\psi_{u}(t_{1};\cdot),\dots,\operatorname{Im}\psi_{u}(t_{M};\cdot)\big)^{\top}.$
Weighted least-squares / GMM criterion.
For a positive semidefinite weighting matrix $W\in\mathbb{R}^{2M\times 2M}$, we estimate $(\upsilon_{u},\sigma_{u}^{2},\Lambda)$ by minimizing
$Q_{u}(\upsilon_{u},\sigma_{u}^{2},\Lambda):=\big(\widehat{Z}_{u}-Z_{u}(\upsilon_{u},\sigma_{u}^{2},\Lambda)\big)^{\top}W\big(\widehat{Z}_{u}-Z_{u}(\upsilon_{u},\sigma_{u}^{2},\Lambda)\big),$ (6)
over $\Lambda$ in the simplex and $\sigma^{2}_{u,j}>0$. Following [76, 82], $W$ may be estimated using HAC-type procedures (e.g. [58]), yielding a generalized method-of-moments estimator in the sense of [42].
Multiple directions and label alignment.
Draw $u_{1},\dots,u_{k}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$ independently and compute, for each $r=1,\dots,k$, the one-dimensional estimate
$(\widehat{\upsilon}_{u_{r},(1)},\widehat{\sigma}^{2}_{u_{r},(1)},\widehat{\Lambda}_{u_{r},(1)}).$
Since each projected fit is performed independently, component labels may be permuted across directions. We align labels as follows.
Choose a pivot direction $u_{r^{\ast}}$ (e.g. the one maximizing the separation between the estimated projected means) and order its components by increasing $\widehat{\upsilon}_{u_{r^{\ast}},(1),j}$. For each $r$, match the $m$ triples $(\widehat{\lambda}_{u_{r},(1),j},\widehat{\upsilon}_{u_{r},(1),j},\widehat{\sigma}^{2}_{u_{r},(1),j})$ to the pivot triples by solving a minimum-cost assignment problem (e.g. Hungarian algorithm), producing aligned estimates $(\widetilde{\upsilon}_{u_{r},(1)},\widetilde{\sigma}^{2}_{u_{r},(1)},\widetilde{\Lambda}_{u_{r},(1)})$. Finally, define the Step 1 weight estimator by averaging:
$\widehat{\Lambda}_{(1)}:=\frac{1}{k}\sum_{r=1}^{k}\widetilde{\Lambda}_{u_{r},(1)}.$ (7)
Remark (Student $t$-mixtures).
For $t$-mixtures, the projected model remains a univariate $t$-mixture with projected locations $\upsilon_{u,j}=\langle u,\mu_{j}\rangle$ and projected scales determined by $\langle u,\Sigma_{j}u\rangle$; Step 1 is identical, replacing $\psi_{u}(\cdot)$ by the corresponding $t$-mixture characteristic function.
A.2 Step 2: Re-estimation of projected means/variances and reconstruction in $\mathbb{R}^{d}$
With $\Lambda$ fixed at $\widehat{\Lambda}_{(1)}$, we re-estimate, for each direction $u_{r}$, the projected location and variance parameters by minimizing the same criterion (6) over $(\upsilon_{u},\sigma_{u}^{2})$ only. Denote the resulting aligned estimators by
$\widehat{\upsilon}_{u_{r},(2)}=(\widehat{\upsilon}_{u_{r},(2),1},\dots,\widehat{\upsilon}_{u_{r},(2),m}),\qquad\widehat{\sigma}^{2}_{u_{r},(2)}=(\widehat{\sigma}^{2}_{u_{r},(2),1},\dots,\widehat{\sigma}^{2}_{u_{r},(2),m}).$
Reconstruction of mean vectors.
For each component $j=1,\dots,m$, we reconstruct $\mu_{j}\in\mathbb{R}^{d}$ via least squares:
$\widehat{\mu}_{j}:=\operatorname*{argmin}_{\mu\in\mathbb{R}^{d}}\ \sum_{r=1}^{k}\big(\langle u_{r},\mu\rangle-\widehat{\upsilon}_{u_{r},(2),j}\big)^{2}.$ (8)
Let $U\in\mathbb{R}^{k\times d}$ be the matrix with $r$-th row $u_{r}^{\top}$ and let $\widehat{v}_{j}=(\widehat{\upsilon}_{u_{1},(2),j},\dots,\widehat{\upsilon}_{u_{k},(2),j})^{\top}$. When $U^{\top}U$ is invertible (which holds with probability one for i.i.d. random directions as soon as $k\geq d$), the solution is
$\widehat{\mu}_{j}=(U^{\top}U)^{-1}U^{\top}\widehat{v}_{j}.$
Reconstruction of covariance matrices.
For each component $j=1,\dots,m$, we reconstruct $\Sigma_{j}\in\mathbb{S}^{+,d}$ by solving the convex quadratic program
$\widehat{\Sigma}_{j}:=\operatorname*{argmin}_{\Sigma\in\mathbb{S}^{+,d}}\ \sum_{r=1}^{k}\big(\langle u_{r},\Sigma u_{r}\rangle-\widehat{\sigma}^{2}_{u_{r},(2),j}\big)^{2}.$ (9)
As discussed in Proposition (Uniqueness of covariance reconstruction) in the main text, uniqueness holds if $\{u_{r}u_{r}^{\top}\}_{r=1}^{k}$ spans $\mathbb{S}^{d}$ (equivalently, if the associated design matrix has rank $d(d+1)/2$). In practice, (9) is solved with a semidefinite-programming routine (e.g. primal–dual interior-point methods, cf. [57, 81, 10]).
Output of Algorithm 1.
The final estimator is
$\widehat{\Theta}=\big(\widehat{\Lambda}_{(1)},\widehat{\bm{\mu}},\widehat{\bm{\Sigma}}\big),\qquad\widehat{\bm{\mu}}=(\widehat{\mu}_{1},\dots,\widehat{\mu}_{m}),\quad\widehat{\bm{\Sigma}}=(\widehat{\Sigma}_{1},\dots,\widehat{\Sigma}_{m}).$
Appendix B Example 3 (Section 3): Constrained-parameter estimation for a $d=20$ three-component Student $t$-mixture
We now consider a high-dimensional Student $t$-mixture in dimension $d=20$, and we compare the estimation obtained by the random projections approach (RP) with a multivariate EM procedure under a restricted parameterization. As in the previous examples, we generate i.i.d. observations from a finite mixture of multivariate $t$ distributions, but here we impose structural constraints on both the component locations and the scatter matrix in order to reduce the dimension of the parameter space.
Specifically, let $t_{\nu}(\mu,\Sigma)$ denote the $d$-variate Student $t$-distribution with $\nu$ degrees of freedom, location parameter $\mu\in\mathbb{R}^{d}$, and (scatter) matrix $\Sigma\in\mathbb{R}^{d\times d}$. We consider the three-component mixture
$F:=\lambda_{1}\,t_{\nu}(\mu_{1},\Sigma)+\lambda_{2}\,t_{\nu}(\mu_{2},\Sigma)+\lambda_{3}\,t_{\nu}(\mu_{3},\Sigma),\qquad\lambda_{1}+\lambda_{2}+\lambda_{3}=1,$
with common degrees of freedom $\nu=2$ and mixing weights
$(\lambda_{1},\lambda_{2},\lambda_{3})=(0.3,0.3,0.4).$
Constrained locations.
To reduce the location-parameter space, we constrain each mean vector to have constant coordinates,
$\mu_{k}=m_{k}\,\mathbf{1}_{d},\qquad k=1,2,3,$
where $\mathbf{1}_{d}$ denotes the $d$-dimensional vector of ones. In the simulation we set
$m_{1}=0,\qquad m_{2}=1,\qquad m_{3}=3,$
so that $\mu_{1}=(0,\ldots,0)$, $\mu_{2}=(1,\ldots,1)$, and $\mu_{3}=(3,\ldots,3)$.
Compound-symmetry covariance.
To reduce the scatter-parameter space, the three components share a common compound-symmetry (CS) matrix,
$\Sigma=\Sigma(x):=(1-x)I_{d}+x\,\mathbf{1}_{d}\mathbf{1}_{d}^{\top},$
that is, $\Sigma_{ii}=1$ and $\Sigma_{ij}=x$ for $i\neq j$. In the simulation we take $x=0.25$ (which must satisfy 1/(d1)<x<1-1/(d-1)<x<1 to ensure positive definiteness).
Sampling scheme.
For each replicate, we generate $N=200$ i.i.d. observations from $F$ by first sampling latent labels $Z_{i}\in\{1,2,3\}$ with $\mathbb{P}(Z_{i}=k)=\lambda_{k}$, and then drawing
$X_{i}\mid(Z_{i}=k)\sim t_{\nu}(\mu_{k},\Sigma),\qquad i=1,\ldots,n.$
This experiment is replicated $100$ times.
Estimation and evaluation.
In each replicate we estimate the mixture parameters under the above constraints using: (i) a multivariate EM algorithm that estimates $(\lambda_{1},\lambda_{2},\lambda_{3})$, $(m_{1},m_{2},m_{3})$ and $x$; and (ii) an RP-based procedure that combines (a) a screened collection of random projections, (b) one-dimensional $t$-mixture fits to recover $(m_{1},m_{2},m_{3})$ and $x$, and (c) a short multivariate refinement initialized at the RP estimates.
For both procedures, we compute MAP allocations and summarize clustering performance via the adjusted Rand index (ARI), see Figure 10. In addition, we measure estimation accuracy through absolute errors for the mixing weights, the mean scalars (reported as errors for $\mu_{1},\mu_{2},\mu_{3}$ under the constraint $\mu_{k}=m_{k}\mathbf{1}_{d}$), see Figures 11, 12, and 13; and the scatter matrix, see Figure 14.
Fig. 10: Boxplots of the Adjusted Rand Index (ARI) over the $100$ Monte Carlo replicates, comparing the EM and RP procedures. Larger ARI values indicate a better agreement between the estimated allocations and the true component labels.
Fig. 11: Boxplots of estimation errors over the $100$ Monte Carlo replicates for component $k=1$: (left) $|\mu_{1}-\mu_{1}^{*}|$ and (right) $|\lambda_{1}-\lambda_{1}^{*}|$, comparing EM and RP. Boxes show the median and interquartile range; whiskers follow Tukey’s rule (outliers omitted in the plot).
Fig. 12: Boxplots of estimation errors over the $100$ Monte Carlo replicates for component $k=2$: (left) $|\mu_{2}-\mu_{2}^{*}|$ and (right) $|\lambda_{2}-\lambda_{2}^{*}|$, comparing EM and RP. Boxes show the median and interquartile range; whiskers follow Tukey’s rule (outliers omitted in the plot).
Fig. 13: Boxplots of estimation errors over the $100$ Monte Carlo replicates for component $k=3$: (left) $|\mu_{3}-\mu_{3}^{*}|$ and (right) $|\lambda_{3}-\lambda_{3}^{*}|$, comparing EM and RP. Boxes show the median and interquartile range; whiskers follow Tukey’s rule (outliers omitted in the plot).
Fig. 14: Boxplots of the Frobenius error $\|\Sigma-\Sigma^{*}\|_{F}$ over the $100$ Monte Carlo replicates, comparing EM and RP.
Appendix C Descriptive Summary Tables and Confusion Matrices for the Simulations in Section 3
Table 1: Average values of the estimated parameters over $100$ replicates of the algorithm. Marginal standard deviations (sd) are shown in parentheses. Results are reported for four separability scenarios indexed by the location-shift parameter $\eta\in\{1/2,1,3/2,2\}$.
Algorithm Component $\mu_{1}$ $\mu_{2}$ $\sigma_{11}$ $\sigma_{22}$ $\sigma_{21}$
Average estimated values
Values 1 0 0 1 1/2 0
2 $\eta$ 0 1/2 1 0
Scenario: $\eta=1/2$
RP 1 -0.36 -0.04 0.92 0.90 0.00
sd (0.94) (1.30) (1.35) (0.65) (0.62)
2 0.92 0.04 0.66 1.02 -0.05
sd (0.65) (1.37) (0.59) (0.88) (0.48)
EM-st 1 0.03 -0.05 0.91 0.63 0.00
sd (1.04) (0.98) (2.61) (0.85) (0.90)
2 0.89 -0.01 0.50 0.64 0.01
sd (0.66) (0.69) (0.48) (0.83) (0.54)
Scenario: $\eta=1$
RP 1 -0.38 0.11 0.84 0.86 0.03
sd (0.71) (1.25) (0.48) (0.49) (0.33)
2 1.35 -0.01 0.57 0.80 0.02
sd (0.49) (1.07) (0.25) (0.45) (0.22)
EM-st 1 0.41 -0.01 0.55 0.50 -0.05
sd (0.91) (0.69) (0.49) (0.40) (0.50)
2 1.48 -0.01 0.71 0.79 -0.28
sd (0.98) (0.76) (2.65) (2.61) (2.65)
Scenario: $\eta=3/2$
RP 1 -0.14 -0.03 1.51 1.10 -0.36
sd (1.00) (1.12) (5.75) (3.47) (4.25)
2 1.73 -0.04 0.66 0.97 0.00
sd (0.45) (1.11) (0.31) (0.99) (0.45)
EM-st 1 0.87 0.13 1.12 0.75 0.17
sd (0.69) (0.97) (3.42) (0.90) (1.59)
2 1.39 0.10 0.44 0.64 -0.02
sd (0.65) (1.05) (0.28) (0.24) (0.30)
Scenario: $\eta=2$
RP 1 -0.27 0.01 1.10 0.65 -0.07
sd (0.39) (0.44) (0.43) (0.50) (0.34)
2 2.20 0.06 0.67 0.89 0.00
sd (0.27) (0.41) (0.28) (0.33) (0.17)
EM-st 1 0.89 -0.03 0.66 0.64 -0.09
sd (0.53) (0.88) (0.39) (0.33) (0.44)
2 1.60 -0.07 0.36 0.74 -0.02
sd (0.49) (1.02) (0.14) (0.26) (0.18)
$\eta$ EM-st RP
1/2 $\begin{pmatrix}0.163&0.139\\ 0.310&0.389\end{pmatrix}$ $\begin{pmatrix}0.133&0.169\\ 0.253&0.446\end{pmatrix}$
1.0 $\begin{pmatrix}0.081&0.188\\ 0.171&0.561\end{pmatrix}$ $\begin{pmatrix}0.150&0.118\\ 0.200&0.532\end{pmatrix}$
3/2 $\begin{pmatrix}0.203&0.093\\ 0.294&0.409\end{pmatrix}$ $\begin{pmatrix}0.198&0.099\\ 0.139&0.564\end{pmatrix}$
2.0 $\begin{pmatrix}0.156&0.141\\ 0.120&0.583\end{pmatrix}$ $\begin{pmatrix}0.206&0.091\\ 0.097&0.607\end{pmatrix}$
Table 2: Average confusion matrices (joint proportions) between true labels and posterior classifications for EM-st and RP.
Table 3: Average values of the estimated parameters over $100$ replicates of the algorithm. Marginal standard deviations (sd) are shown in parentheses. Results are reported for three contamination scenarios indexed by the outlier proportion $\gamma\in\{0.05,0.10,0.15\}$.
Algorithm Component $\mu_{1}$ $\mu_{2}$ $\sigma_{11}$ $\sigma_{22}$ $\sigma_{21}$
Average estimated values
Values 1 0 0 1 1/2 0
2 2 0 1/2 1 0
Scenario: $\gamma=0.05$
RP 1 -0.27 0.12 1.16 0.68 0.00
sd (0.32) (0.54) (0.41) (0.53) (0.31)
2 2.27 0.04 0.72 1.05 0.05
sd (0.40) (0.43) (0.45) (0.31) (0.12)
EM-rob 1 0.60 0.10 3.18 1.32 0.17
sd (0.32) (0.17) (0.75) (0.39) (0.43)
2 2.07 0.05 0.89 2.31 0.06
sd (0.10) (0.17) (0.37) (0.51) (0.16)
Scenario: $\gamma=0.10$
RP 1 -0.14 0.46 1.08 0.73 0.07
sd (0.45) (0.86) (0.27) (0.40) (0.20)
2 2.24 -0.11 0.68 1.09 0.05
sd (0.37) (0.68) (0.18) (0.40) (0.14)
EM-rob 1 0.70 0.17 3.33 1.61 0.34
sd (0.29) (0.23) (0.68) (0.66) (0.70)
2 2.07 0.13 0.93 2.53 0.06
sd (0.10) (0.23) (0.36) (0.74) (0.18)
Scenario: $\gamma=0.15$
RP 1 0.12 0.59 1.03 0.88 0.09
sd (0.63) (1.75) (0.35) (0.40) (0.24)
2 2.21 0.28 0.70 1.12 0.09
sd (0.42) (1.57) (0.25) (0.81) (0.21)
EM-rob 1 0.77 0.28 3.37 1.85 0.54
sd (0.31) (0.30) (0.73) (0.87) (0.84)
2 2.08 0.21 1.07 2.80 0.08
sd (0.11) (0.29) (0.36) (1.02) (0.22)
$\gamma$ RobEM RP
$\gamma=5\%$ $\begin{pmatrix}0.262&0.043\\ 0.157&0.538\end{pmatrix}$ $\begin{pmatrix}0.232&0.072\\ 0.063&0.632\end{pmatrix}$
$\gamma=10\%$ $\begin{pmatrix}0.254&0.046\\ 0.140&0.560\end{pmatrix}$ $\begin{pmatrix}0.191&0.109\\ 0.064&0.636\end{pmatrix}$
$\gamma=15\%$ $\begin{pmatrix}0.253&0.045\\ 0.178&0.524\end{pmatrix}$ $\begin{pmatrix}0.292&0.076\\ 0.103&0.529\end{pmatrix}$
Table 4: Average confusion matrices (joint proportions) between true labels and posterior classifications for RobEM-st and RP.